### v1

# The Dirichlet problem for hessian quotient type curvature equations in Minkowski space

Abstract.
In this paper, we consider the Dirichlet problem for a class of prescribed Hessian quotient type curvature equations in Minkowski space. By establishing the a priori $C^{2}$ estimates, we prove the existence of smooth spacelike hypersurfaces with a class of prescribed Hessian quotient type curvature and homogeneous boundary data in the non-degenerate case.
Keywords: Minkowski space; Hessian quotient type curvature equations; a priori $C^{2}$ estimates.
1. Introduction
In this paper, we consider the Dirichlet problem for the following non-degenerate Hessian quotient type curvature equations in Minkowski space $\mathbb{R}^{n,1}$:
(1.1) $\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))=\psi(X).$
When $k=n$ and $l=0$, this problem was studied by H. Jiao and the first author in [15] for general boundary data.
In Euclidean space, Chen-Tu-Xiang [9] derived the existence and uniqueness results for the Dirichlet problem of
$\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))=\psi(X,\nu),\;\;k<n$
with homogeneous boundary data in degenerate form. The current work aims to extend the results of [9] from Euclidean space to Minkowski space $\mathbb{R}^{n,1}$, whose induced metric is given by
$ds^{2}=dx^{2}_{1}+\cdots+dx^{2}_{n}-dx^{2}_{n+1}.$
Due to the lack of gradient estimates, we restrict our analysis to the non-degenerate case $\psi=\psi(X)>0$, see Remark 3.2.
In contrast to Euclidean space, when differentiating equations twice in the Minkowski space, uncontrollable terms including the square of curvatures will appear. This is the reason why the Dirichlet problem for $k$-curvature equations
(1.2) $\sigma_{k}(\kappa[M_{u}])=\psi$
remains unsolved for $33\leqslant k\leqslant n-$ 3, although it has been solved a long time ago for corresponding equations in Euclidean space, where
$\sigma_{k}(\kappa[M_{u}])=\sum_{1\leqslant i_{1}<\cdots<i_{k}\leqslant n}\kappa_{i_{1}}\cdots\kappa_{i_{k}},$
where $\kappa[M_{u}]=(\kappa_{1},\cdots,\kappa_{n})$ denote the principal curvatures of $M_{u}$. The reader is referred to [4, 16, 17, 22, 18, 6, 7, 30] for the study of the Dirichlet problem for (1.2) and the quotient type curvature equation in Euclidean space.
In this paper, we are interested in the graph $M_{u}$ of a spacelike function $u$ defined on a bounded domain $\Omega\subset\mathbb{R}^{n}$, which is
$M_{u}:=\{(x,u(x)):x\in\Omega\subset\mathbb{R}^{n}\}.$
For the function $v(x)\in C^{2}(\Omega)$, we denote $v_{i}=\frac{\partial v}{\partial x_{i}}$, $v_{ij}=\frac{\partial^{2}v}{\partial x_{i}\partial x_{j}}$, and $Dv=(v_{1},\cdots,v_{n})$ and $D^{2}v=(v_{ij})$ represent the ordinary gradient and Hessian matrix.
In Minkowski space, we call a function $u$ spacelike if
$\sup_{\overline{\Omega}}|Du|<1.$
Clearly, for the spacelike function $u$, the Minkowski metric restricted to $M_{u}$ induces a Riemannian metric on $M_{u}$:
[…]
The inverse of the metric is
[…]
The second fundamental form $M_{u}$ with respect to its upward unit normal vector field
$\nu=\frac{(Du,1)}{\sqrt{1-|Du|^{2}}}$
is given by
$h_{ij}=\frac{u_{ij}}{\sqrt{1-|Du|^{2}}}.$
Define the $(0,2)$-tensor field $\eta$ on $M$ by
$\eta=Hg-h,$
where $H(X)$ is the mean curvature of $M_{u}$ at $X=(x,u(x))\in M_{u}$, $g=\{g_{ij}\}$, $h=\{h_{ij}\}$. $\lambda(\eta[M_{u}])=(\lambda_{1},\cdots,\lambda_{n})$ denote the eigenvalues of $\eta$ with respect to $g$ on $M_{u}$.
In this paper, we consider the existence of smooth spacelike graphic hypersurface satisfying the Dirichlet problem
(1.3) $\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))&=\psi(x,u)&\;\;~{}\mbox{ in }\Omega,\\ u&=0&\;\;~{}\mbox{ on }\partial\Omega,\end{aligned}\right.$
where $\psi\in C^{\infty}(\overline{\Omega}\times\mathbb{R})>0$.
We first introduce the definition of admissible set of equation (1.1).
Definition 1.1.
A $C^{2}$ regular spacelike hypersurface $M_{u}$ is called $(\eta,k)$-convex if the principal curvatures $\kappa=(\kappa_{1},\ldots,\kappa_{n})\in\widetilde{\Gamma}_{k}$ at each $X\in M_{u}$, where $\widetilde{\Gamma}_{k}$ is the symmetric cone defined by
(1.4) $\widetilde{\Gamma}_{k}:=\{\kappa=(\kappa_{1},\ldots,\kappa_{n})\in\mathbb{R}^{n}:\lambda(\eta)=(\lambda_{1},\cdots,\lambda_{n})\in\Gamma_{k},\lambda_{i}(\eta)=\sum_{j\neq i}\kappa_{j}\}.$
$\Gamma_{k}$ is defined by
$\Gamma_{k}:=\{\lambda=(\lambda_{1},\cdots,\lambda_{n})\in\mathbb{R}^{n}:\sigma_{m}(\lambda)>0,m=1,\cdots,k\}.$
In addition, we call a $C^{2}$ spacelike function $u:\Omega\rightarrow\mathbb{R}$ admissible if its graph $M_{u}$ is $(\eta,k)$-convex.
We also need some geometric conditions on $\Omega$ similar to those in [17] when consider the Dirichlet boundary value problem.
Definition 1.2.
If $k<n$, a bounded domain $\Omega\subset\mathbb{R}^{n}$ is called $(\eta,k)$-convex, if there exists a positive constant $K$ such that for each $x\in\partial\Omega$,
$(\kappa^{b}_{1}(x),\ldots,\kappa^{b}_{n-1}(x),K)\in\widetilde{\Gamma}_{k},$
where $\kappa^{b}_{1}(x),\ldots,\kappa^{b}_{n-1}(x)$ are the principal curvatures of $\partial\Omega$ at $x$.
Jiao and Sun [20] considered the following degenerate curvature equation
$\det(\lambda(\eta[M_{u}]))=\psi(X,\nu)$
in Euclidean space with constant boundary condition. They derived the existence of $C^{1,1}$ regular graphical $(\eta,n)$-convex hypersurfaces. Motivated by [20], Chen-Tu-Xiang [9] extended this work to a degenerate Hessian quotient type curvature equation (1.1) with $0\leqslant l<k<n$, they proved the existence and uniqueness of $C^{1,1}$ regular graphical hypersurfaces in Euclidean space under the condition $\psi^{\frac{1}{k-1}}(x,u(x),\nu(x))\in C^{1,1}(\Omega\times\mathbb{R}\times S^{n})$. Building on [9], we generalize their results to Minkowski space and establish the following theorem:
Theorem 1.3.
Let $k\geqslant 2,~{}0\leqslant l<k<n$. Assume that bounded domain $\Omega\subset\mathbb{R}^{n}$ is $(\eta,k)$-convex with smooth boundary $\partial\Omega$. Suppose $\psi=\psi(x,z)\in C^{\infty}(\overline{\Omega}\times\mathbb{R})>0$, $\psi_{z}\geqslant 0$. In addition, assume that there exists an admissible subsolution $\underline{u}\in C^{2}(\overline{\Omega})$ satisfying $\kappa[M_{\underline{u}}]\in\widetilde{\Gamma}_{k}$ and
[…]
Then there exists a unique admissible solution $u\in C^{\infty}(\overline{\Omega})$ to (1.3).
To conclude the introduction, we briefly mention some related studies. If $\lambda(\eta[M_{u}])$ is replaced by the principal curvature $\kappa[M_{u}]$ of $M_{u}$ and $l=0$, equation (1.1) becomes the classical $k$-curvature equations (1.2). The mean curvature, scalar curvature, and Gauss curvature correspond respectively to $k=1,2,n$ in the formula (1.2).
The Dirichlet problem for prescribed mean curvature in Minkowski space was first solved by Bartnik and Simon [1], while Delanoè [10] studied the Dirichlet problem of the prescribed Gauss curvature equation (see also [11]). Bayard [2] completely solved the solvability of the Dirichlet problem of the prescribed scalar curvature equation in four-dimensional space under the condition of strict convexity, and Urbas [28] generalized the result to the general dimension $n$. Ren and Wang [24, 25] solved the case $k=n-1$ and $k=n-2$. For $33\leqslant k\leqslant n-$ 3, this is still an open problem. The reason is the lack of a priori $C^{2}$ estimates. For this problem, Guan-Ren-Wang [14] given a second order estimates for admissible solutions of $k$-curvature equations (1.2) whose graphs are $(k+1)$-convex, and Wang [29] established the estimates for admissible solutions whose graphs are semi-convex.
For the corresponding Hessian quotient type equations
$\frac{\sigma_{k}(\lambda(\Delta uI-D^{2}u))}{\sigma_{l}(\lambda(\Delta uI-D^{2}u))}=f(x,u(x),Du(x)),$
Liu-Mao-Zhao [23] established Pogorelov type estimates of $k$-convex solutions with the vanishing Dirichlet boundary condition in Lorentz-Minkowski space $\mathbb{R}^{n+1}$. Chen-Tu-Xiang [8] considered the Dirichlet problem for the above equation on compact Riemannian Manifolds and obtained the existence and uniqueness.
The rest of this paper is organized as follows. In Section 2, we provided some preliminaries. In Section 3, The $C^{1}$ estimates are established. Section 4 is devoted to estimates for second order derivatives.
2. Preliminaries
Let $\epsilon_{1},\cdots,\epsilon_{n+1}$ be an orthonormal basis of $\mathbb{R}^{n+1}$. Consider a spacelike function $u$ and let $M_{u}$ represent its graph, which can be written as
$M_{u}=\{X=(x,u(x))|x\in\mathbb{R}^{n}\}.$
[…]
On the other hand, when we choose local orthonormal frames $\{e_{1},e_{2},\cdots,e_{n}\}$ on $TM_{u}$. $\nabla$ denotes the induced Levi-Civita connection on $M$. For a smooth function $v$ on $M_{u}$, we denote $\nabla_{i}v=\nabla_{e_{i}}v,$ $\nabla_{ij}v=\nabla^{2}v(e_{i},e_{j}),$ etc in this paper. Thus, we have
$|\nabla u|=\sqrt{g^{ij}u_{i}u_{j}}=\frac{|Du|}{\sqrt{1-|Du|^{2}}}.$
Under the orthonormal coordinates $\{e_{1},e_{2},\cdots,e_{n}\}$ introduced earlier, we also have the following fundamental formulae and equations for the hypersurface $M$ in Minkowski space $\mathbb{R}^{n,1}$:
(2.1) $\nabla_{ij}X=$ $h_{ij}\nu\quad{\rm(Gauss\ formula)}$
$\nabla_{i}\nu=$ $h_{ij}e_{j}\quad{\rm(Weigarten\ formula)}$
$\nabla_{k}h_{ij}=$ $\nabla_{j}h_{ik}\quad{\rm(Codazzi\ equation)}$
$R_{ijst}=$ $-(h_{is}h_{jt}-h_{it}h_{js})\quad{\rm(Gauss\ equation)},$
where $h_{ij}=\langle D_{e_{i}}\nu,e_{j}\rangle$ is the second fundamental form of $M$.
Define the martix $A=\{a_{ij}\}$, where
(2.2) $a_{ij}=\frac{1}{w}\gamma^{ik}u_{kl}\gamma^{lj},$
and $w=\sqrt{1-|Du|^{2}}$, $\gamma^{ik}=\delta_{ik}+\frac{u_{i}u_{k}}{w(1+w)}.$ Note that $\{\gamma^{ij}\}$ is invertible and its inverse is $\{\gamma_{ij}\}=\{\delta_{ij}-\frac{u_{i}u_{j}}{1+w}\}$, which is the square root of $\{g_{ij}\}.$
Clearly, $A=\{a_{ij}\}$ is symmetric and the eigenvalues of $A$ equal the prescribed principal curvatures of $M_{u}$, which means
$\lambda(\{a_{ij}\})=\lambda(\{g^{ij}\}\{h_{ij}\})=\lambda\big{(}\frac{1}{w}\left(I+\frac{Du\otimes Du}{w^{2}}\right)D^{2}u\big{)}.$
Note that the equation (1.1) can be rewritten by
$f(\kappa):=\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))=\psi,$
where $\kappa=(\kappa_{1},\ldots,\kappa_{n})$ are the principal curvatures of $M_{u}$. We need the following properties of $f(\kappa)$ (seeing [19]). For any constant $A>0$ and any compact set $K$ in $\widetilde{\Gamma}_{k}$ there is a number $R=R(A,K)$ such that
(2.3) $f(\kappa_{1},\ldots,\kappa_{n-1},\kappa_{n}+R)\geqslant A,\mbox{ for all }\kappa\in K.$
In the final part of this section, we list some algebraic equalities and inequalities of $\sigma_{k}$. Denote $\sigma_{m;i_{1},\cdots,i_{k}}(\kappa)=\sigma_{m}(\kappa)|_{\kappa_{i_{1}}=\cdots=\kappa_{i_{k}}=0}$ for integer $1\leqslant i_{1},\cdots,i_{k}\leqslant n,1\leqslant m\leqslant n$ and $n-k\leqslant m$. Then we have:
$\frac{\partial\sigma_{k}}{\partial\kappa_{i}}(\kappa)=\sigma_{k-1;i}(\kappa),$
$\sum_{i}\sigma_{k-1;i}(\kappa)=(n-k+1)\sigma_{k-1}(\kappa)$
and
$\sum_{i}\sigma_{k-1;i}(\kappa)\kappa_{i}=k\sigma_{k}(\kappa).$
The generalized Newton-MacLaurin inequality is as follows.
Proposition 2.1.
For $\kappa\in\Gamma_{k}$ and $n\geqslant k>l\geqslant 0$, $n\geqslant r>s\geqslant 0$, $k\geqslant r$, $l\geqslant s$, we have
$\Bigg{[}\frac{{\sigma_{k}(\kappa)}/{C_{n}^{k}}}{{\sigma_{l}(\kappa)}/{C_{n}^{l}}}\Bigg{]}^{\frac{1}{k-l}}\leqslant\Bigg{[}\frac{{\sigma_{r}(\kappa)}/{C_{n}^{r}}}{{\sigma_{s}(\kappa)}/{C_{n}^{s}}}\Bigg{]}^{\frac{1}{r-s}}.$
Proof.
See [26]. ∎
Next, we introduce a new cone $\widetilde{\Gamma}_{k}$ defined by (1.4) and list some important properties, more details can refer to [5].
Proposition 2.2.
The following properties hold.
(1) $\widetilde{\Gamma}_{k}$ are convex cones and
(2.4) $\Gamma_{1}=\widetilde{\Gamma}_{1}\supset\widetilde{\Gamma}_{2}\supset\cdots\supset\widetilde{\Gamma}_{n}\supset\Gamma_{2}.$
(2) If $\kappa=(\kappa_{1},\cdots,\kappa_{n})\in\widetilde{\Gamma}_{k}$, then $\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]^{\frac{1}{k-l}}(0\leqslant l<k\leqslant n)$ are concave with respect to $\kappa$. Hence for any $(\xi_{1},\cdots,\xi_{n})$, we have
(2.5) $\sum_{i,j}\frac{\partial^{2}\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\kappa_{i}\partial\kappa_{j}}\xi_{i}\xi_{j}\leqslant\left(1-\frac{1}{k-l}\right)\frac{\left[\sum_{i}\frac{\partial\left(\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right)}{\partial\kappa_{i}}\xi_{i}\right]^{2}}{\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}}.$
Proposition 2.3.
Suppose that $\kappa=(\kappa_{1},\cdots,\kappa_{n})\in\widetilde{\Gamma}_{k}$ are ordered with $\kappa_{1}\geqslant\kappa_{2}\geqslant\cdots\geqslant\kappa_{n}$, then for $0\leqslant l<k<n$, we have
(2.6) $\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\kappa_{i}}\geqslant c(n,k,l)\sum_{i}\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\kappa_{i}}\geqslant c(n,k,l)\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]^{1-\frac{1}{k-l}},~{}\forall~{}1\leqslant i\leqslant n.$
3. $C^{1}$ estimates
In this section, we establish the $C^{1}$ estimates for the admissible solution to (1.3). Simliar as [15], we prove the admissible solution $u$ is spacelike, which means there exists a constant $10<\theta_{0}<$ 1 such that
(3.1) $\sup_{\overline{\Omega}}|Du|\leqslant 1-\theta_{0}.$
Since $M_{u}$ is $(\eta,k)$-convex, then $\sigma_{1}(\lambda(\eta[M_{u}]))>0$, hence $H[M_{u}]>0$. By proposition (2.1), we can find a spacelike supersolution $\overline{u}\in C^{\infty}(\overline{\Omega})$ of (1.3). Then combining with $\psi_{u}\geqslant 0$, the subsolution condition (1.5) and the maximum principle, we can find that
(3.2) $\sup_{\overline{\Omega}}|u|\leqslant\max\{\sup_{\overline{\Omega}}|\overline{u}|,\sup_{\overline{\Omega}}|\underline{u}|\}\leqslant C$
and
(3.3) $\sup_{\partial\Omega}|Du|\leqslant\max\{\sup_{\partial\Omega}|D\underline{u}|,\sup_{\partial\Omega}|D\overline{u}|\}\leqslant 1-\theta$
for some constant $10<\theta<$ 1 since $\underline{u}$ and $\overline{u}$ are both spacelike. Next we establish the global gradient estimates (3.1), that is, we prove that the function
$\tilde{w}=\frac{1}{\sqrt{1-|Du|^{2}}}=\frac{1}{w}.$
has an upper bound.
Theorem 3.1.
Let $u\in C^{3}(\Omega)\cap C^{1}(\overline{\Omega})$ be an admissible solution of (1.3). Suppose the smooth function $\psi$ satisfies $\psi_{u}\geqslant 0$. Then
(3.4) $\sup_{\overline{\Omega}}\tilde{w}\leqslant C(1+\sup_{\partial\Omega}\tilde{w}),$
where $C$ depends on $|D\psi|$, the lower bound of $\psi$, $\sup_{\overline{\Omega}}|u|$ and other known data.
Proof.
Take barrier function
$\hat{Q}:=\tilde{w}e^{Bu},$
where $B$ is a positive constant to be determined later. Suppose the maximum value of $\hat{Q}$ is achieved at an interior point $x_{0}\in\Omega$. Then
$Q:=\log\hat{Q}=\log\tilde{w}+Bu$
also attains its maximum at $x_{0}$. Let $\epsilon_{1},\ldots,\epsilon_{n+1}$ be a standard basis of $\mathbb{R}^{n+1}$. We rotate the coordinate system $\epsilon_{1},\ldots,\epsilon_{n}$ so that $u_{1}(x_{0})=|Du(x_{0})|$ and $u_{j}(x_{0})=0$ for $j\geqslant 2$. Next, we consider on the manifold $M_{u}$. Define
$e_{i}=\gamma^{is}\tilde{\partial}_{s},\ i=1,\ldots,n,$
where $\gamma^{is}:=\delta_{is}+\frac{u_{i}u_{s}}{w(1+w)}$ and ~$\tilde{\partial}_{s}:=\epsilon_{s}+u_{s}\epsilon_{n+1}$. It’s clear that $\{e_{1},e_{2},\cdots,e_{n}\}$ be an orthonormal frame on $TM_{u}$ around $X_{0}=(x_{0},u(x_{0}))$ such that, at $x_{0}$,
(3.5) $\nabla_{1}u=\frac{|Du|}{w}=|\nabla u|,\nabla_{i}u=u_{i}=0,\text{ for }i\geqslant 2.$
We may further rotate $\epsilon_{2},\ldots,\epsilon_{n}$ such that $\{u_{ij}\}_{i,j\geqslant 2}$ is diagonal at $x_{0}$. At $x_{0}$, we have
$0=\nabla_{i}Q=\frac{\nabla_{i}\tilde{w}}{\tilde{w}}+B\nabla_{i}u$
Note that by the Weingarten formula,
$\nabla_{i}\tilde{w}=-\nabla_{i}\langle\nu,\epsilon_{n+1}\rangle=-\langle h_{ij}e_{j},\epsilon_{n+1}\rangle=-h_{ij}\langle\gamma^{js}\tilde{\partial_{s}},\epsilon_{n+1}\rangle$
and
$\langle\gamma^{js}\tilde{\partial_{s}},\epsilon_{n+1}\rangle=-\gamma^{js}u_{s}=-\frac{u_{j}}{w}=-\nabla_{j}u.$
Combing (3.5), we have, at $x_{0}$
(3.6) $0=\nabla_{i}Q=\frac{h_{i1}\nabla_{1}u}{\tilde{w}}+B\nabla_{i}u$
and
(3.7) $0\geqslant F^{ii}\nabla_{ii}Q=\frac{F^{ii}\nabla_{i}h_{i1}\nabla_{1}u+F^{ii}h_{il}\nabla_{il}u}{\tilde{w}}-\frac{F^{ii}\left(h_{i1}\nabla_{1}u\right)^{2}}{\tilde{w}^{2}}+BF^{ii}\nabla_{ii}u,$
where
$F^{ij}:=\frac{\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta))}{\partial h_{ij}}.$
In the rest of the proof, we declare that all computations are carried out at $X_{0}$. From (3.6), we have
$h_{11}=-B\tilde{w}\mbox{ and }h_{1i}=0\mbox{ for }i\geqslant 2.$
Since at $X_{0}$,
$h_{11}=\frac{u_{11}}{w^{3}},h_{1i}=\frac{u_{1i}}{w^{2}}\mbox{ and }h_{ij}=\frac{u_{ij}}{w}\mbox{ for }i,j\geqslant 2,$
we find that the matrix $\{h_{ij}\}$ is diagonal at $X_{0}$, and so is $\{F^{ij}\}$. By the Codazzi equation and differentiating the equation (1.1), we have
(3.8) $F^{ii}\nabla_{i}h_{i1}=$ $F^{ii}\nabla_{1}h_{ii}=\nabla_{1}\psi=\psi_{x_{j}}\nabla_{1}x_{j}+\psi_{u}\nabla_{1}u$
$=$ $\frac{\psi_{x_{1}}}{w}+\psi_{u}\nabla_{1}u=\tilde{w}(\psi_{x_{1}}+\psi_{u}u_{1})\geqslant\tilde{w}{\psi_{x_{1}}},$
where the last inequality is due to that $\psi_{u}\geqslant 0$. By the Gauss formula, we have
(3.9) $\nabla_{ij}u=-\nabla_{ij}\langle X,\epsilon_{n+1}\rangle=-\langle\nabla_{ij}X,\epsilon_{n+1}\rangle=-h_{ij}\langle\nu,\epsilon_{n+1}\rangle=\tilde{w}h_{ij}.$
Combining (3.5)-(3.9) and (2.6), we obtain
(3.10) $0\geqslant$ $\psi_{x_{1}}u_{1}\tilde{w}+F^{ii}h_{ii}^{2}-B^{2}F^{11}(\nabla_{1}u)^{2}+B(k-l)\psi\tilde{w}$
$\geqslant$ $-|D\psi|\cdot|Du|\tilde{w}+B^{2}\tilde{w}^{2}F^{11}-B^{2}F^{11}(\nabla_{1}u)^{2}+B(k-l)\psi\tilde{w}$
$\geqslant$ $-|D\psi|\tilde{w}+B^{2}F^{11}+B(k-l)\psi\tilde{w}$
$\geqslant$ $(B(k-l)\psi-|D\psi|)\tilde{w}.$
Then we get a contradiction provided
$B>\frac{\sup_{\overline{\Omega}}|D\psi|}{(k-l)\inf_{\overline{\Omega}}\psi}$
and (3.4) follows as [2]. It follows from (3.3) and (3.4) that (3.1) holds. ∎
Remark 3.2.
Although Chen-Tu-Xiang [9] has given the gradient estimate for admissible solutions of the equation
$\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))=\psi(x,u,\nu)\geqslant 0$
in Euclidean space, gradient estimates in Minkowski space remain a challenge for degenerate problems due to the metric. In addition, no satisfactory approach has been developed to control the term $\nu$, even in non-degenerate cases.
4. $C^{2}$ estimates
We shall first estimate the second derivatives of $u$ on the boundary.
Theorem 4.1.
Suppose bounded domain $\Omega\subset\mathbb{R}^{n}$ is $(\eta,k)$-convex with smooth boundary $\partial\Omega$. Let $u\in C^{3}(\overline{\Omega})$ be an admissible solution of (1.3). Then there exists a positive constant $C$ depending only on $\theta_{0}$, $\|\psi\|_{C^{1}(\overline{\Omega}\times[-\mu_{0},\mu_{0}])}$, $\|\varphi\|_{C^{3}(\overline{\Omega})}$ and $\partial\Omega$ satisfying
(4.1) $\max_{\partial\Omega}|D^{2}u|\leqslant C,$
where $\mu_{0}:=\|u\|_{C^{0}(\overline{\Omega})}$.
We will divide this estimate (4.1) into three parts for proof: the double tangential derivative estimate, the mixed tangential-normal derivative estimate, and the double normal derivative estimate on $\partial\Omega$.
For any point $x_{0}\in\partial\Omega$, without loss of generality, we may assume that $x_{0}$ is the origin and that the positive $x_{n}$-axis is the inner normal direction to $\partial\Omega$ at the origin. Furthermore, we may suppose that in a neighbourhood of the origin, the boundary $\partial\Omega$ is given by
(4.2) $x_{n}=\rho(x^{\prime})=\frac{1}{2}\sum_{\alpha<n}\kappa^{b}_{\alpha}x_{\alpha}^{2}+O(|x^{\prime}|^{3}),$
where $\kappa^{b}_{1},\ldots,\kappa^{b}_{n-1}$ are the principal curvatures of $\partial\Omega$ at the origin and $x^{\prime}=(x_{1},\ldots,x_{n-1})$. Define
$\omega_{\delta}=\{x\in\Omega:\rho(x^{\prime})<x_{n}<\rho(x^{\prime})+\delta^{2},|x^{\prime}|<\delta\},$
we can find that the boundary $\partial\omega_{\delta}$ consists of three parts:
$\partial\omega_{\delta}=\partial_{1}\omega_{\delta}\cup\partial_{2}\omega_{\delta}\cup\partial_{3}\omega_{\delta},$
where $\partial_{1}\omega_{\delta}$, $\partial_{2}\omega_{\delta}$ and $\partial_{3}\omega_{\delta}$ are defined by $\{x_{n}=\rho\}\cap\overline{\omega}_{\delta}$, $\{x_{n}=\rho+\delta^{2}\}\cap\overline{\omega}_{\delta}$ and $\{|x^{\prime}|=\delta\}\cap\overline{\omega}_{\delta}$ respectively.
Part 1. the double tangential derivative estimate on $\partial\Omega$, that is
(4.3) $|u_{\alpha\beta}(0)|\leqslant C\mbox{ for }1\leqslant\alpha,\beta\leqslant n-1,$
Since (4.2), We can obtain (4.3) by differentiating the boundary condition $u=0$ twice on $\partial\Omega$.
Part 2. the mixed tangential-normal derivative estimate on $\partial\Omega$, that is
(4.4) $|u_{\alpha n}(0)|\leqslant C\mbox{ for }1\leqslant\alpha\leqslant n-1.$
We rewrite the equation (1.1) by the form
(4.5) $G(D^{2}u,Du):=\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta(M_{u})))=f(\lambda(A[u]))=\psi(x,u),$
where $G=G(r,p)$ is viewed as a function of $(r,p)$ for $r\in S^{n\times n}$ and $p\in\mathbb{R}^{n}$. Define
(4.6) $G^{ij}=\frac{\partial G}{\partial r_{ij}}(D^{2}u,Du),\ \ G^{i}=\frac{\partial G}{\partial p_{i}}(D^{2}u,Du)$
and the linearized operator by
$L=G^{ij}\partial_{ij}.$
We need the following lemma:
Lemma 4.2.
We have
(4.7) $G^{s}=\frac{u_{s}}{w^{2}}\sum_{i}f_{i}\kappa_{i}+\frac{2}{w(1+w)}\sum_{t,j}F^{ij}a_{it}\big{(}wu_{t}\gamma^{sj}+u_{j}\gamma^{ts}\big{)},$
where $w=\sqrt{1-|Du|^{2}}$, $a_{ij}=\frac{1}{w}\gamma^{ik}u_{kl}\gamma^{lj}$, $\kappa=\lambda(\{a_{ij}\})$, $f_{i}=\frac{\partial f(\kappa)}{\kappa_{i}}$ and
$F^{ij}=\frac{\partial f(\lambda(A[u]))}{\partial a_{ij}}.$
The details of the proof can be found in [15].
Besides, to prove (4.4), we shall use the strategy of [16] to consider the function
$W:=\nabla^{\prime}_{\alpha}u-\frac{1}{2}\sum_{1\leqslant\beta\leqslant n-1}u_{\beta}^{2}$
defined on $\overline{\omega}_{\delta}$ for small $\delta$, where
$\nabla^{\prime}_{\alpha}u:=u_{\alpha}+\rho_{\alpha}u_{n},\mbox{ for }1\leqslant\alpha\leqslant n-1.$
Then we have
$W_{s}=u_{\alpha s}+\rho_{\alpha s}u_{n}+\rho_{\alpha}u_{ns}-\sum_{\beta=1}^{n-1}u_{\beta}u_{\beta s}$
and
$W_{ij}=$ $u_{\alpha ij}+\rho_{\alpha ij}u_{n}+\rho_{\alpha i}u_{nj}+\rho_{\alpha j}u_{ni}$
$+\rho_{\alpha}u_{nij}-\sum_{\beta=1}^{n-1}(u_{\beta i}u_{\beta j}+u_{\beta}u_{\beta ij}).$
We also need the following lemma:
Lemma 4.3.
If $\delta$ is sufficiently small, we have
(4.8) $LW\leqslant C\left(1+|DW|+\sum_{i}G^{ii}+G^{ij}W_{i}W_{j}\right),$
where $C$ is a positive constant depending on $n$, $\theta_{0}$, $\|\psi\|_{C^{1}(\overline{\Omega}\times[-\mu_{0},\mu_{0}]}$ and $\partial\Omega$, where $\mu_{0}=\|u\|_{C^{0}(\overline{\Omega})}$.
Assume (4.8) is correct, next, we establish the mixed tangential-normal derivative estimate (4.4).
Since $\Omega$ is $(\eta,k)$-convex, there exist two positive constants $\theta$ and $K$ satisfying
(4.9) $(\kappa_{1}^{b}-3\theta,\ldots,\kappa_{n-1}^{b}-3\theta,2K)\in\widetilde{\Gamma}_{k}.$
Define
(4.10) $v=\rho(x^{\prime})-x_{n}-\theta|x^{\prime}|^{2}+Kx_{n}^{2}$
on $\overline{\omega}_{\delta}$. We see that when $\delta$ depending on $\theta$ and $K$ is sufficiently small, we have
[…]
In view of (4.2) and (4.9), $\lambda(D^{2}v)\in\widetilde{\Gamma}_{k}$ on $\overline{\omega}_{\delta}$. Thus, there exists an uniform constant $\eta_{0}>0$ depending only on $\theta$, $\partial\Omega$ and $K$ satisfying
$\lambda(D^{2}v-2\eta_{0}I)\in\widetilde{\Gamma}_{k}\mbox{ on }\overline{\omega}_{\delta}.$
Then by a similar calculation of Lemma 3.1 in [17] or Proposition 2.1 in [21], we have
(4.12) $\lambda\left(\frac{1}{w}\{\gamma^{is}(v_{st}-2\eta_{0}\delta_{st})\gamma^{jt}\}\right)\in\widetilde{\Gamma}_{k}\mbox{ on }\overline{\omega}_{\delta}.$
Actually, we can assume $D^{2}v=\operatorname{diag}(\kappa_{1}^{b}-2\theta,\dots,\kappa_{n-1}^{b}-2\theta,2K)$ by rotating the coordinate system if necessary and note that $\gamma^{ij}=\operatorname{diag}(1,\dots,1,\frac{1}{w})$ at 00 since $|Du(0)|=u_{n}(0)$. It follows that
$\{\frac{1}{w}(\gamma^{ik}D_{kl}v\gamma^{lj})\}=\operatorname{diag}(\frac{\kappa_{1}^{b}-2\theta}{w},\dots,\frac{\kappa_{n-1}^{b}-2\theta}{w},\frac{2K}{w^{3}}).$
Let $\kappa_{v}=(\frac{\kappa_{1}^{b}-2\theta}{w},\dots,\frac{\kappa_{n-1}^{b}-2\theta}{w},\frac{2K}{w^{3}})=(\kappa_{v}^{\prime},\frac{2K}{w^{3}})$, we have
$\frac{\sigma_{k}}{\sigma_{l}}(\kappa_{v})>\frac{\sigma_{k}}{\sigma_{l}}(\kappa_{v}^{\prime},\frac{2K}{w})>0$
by $w<1$, (4.9) and (2.6). Then we can find a suitable constant $\delta$ such that (4.12) holds.
As [21] and [20], we consider the following barrier on $\overline{\omega}_{\delta}$, for sufficiently small $\delta$,
(4.13) $\Psi:=v-td+\frac{N}{2}d^{2},$
where $v(x)$ is defined by (4.10), $d(x):=\mathrm{dist}(x,\partial\Omega)$ is the distance from $x$ to the boundary $\partial\Omega$, and $t,N$ are two positive constants to be determined later. Since $f^{1/(k-l)}$ is concave in $\Gamma$ and homogeneous of degree one and $|Dd|\equiv 1$ on the boundary $\partial\Omega$, by (4.12) and (2.3), we have,
$\frac{1}{k-1}\psi^{\frac{1}{k-l}-1}G^{ij}(D^{2}v-\eta_{0}I+NDd\otimes Dd)_{ij}$
$\geqslant$ $G^{1/{(k-l)}}(D^{2}v-\eta_{0}I+NDd\otimes Dd,Du)$
$\geqslant$ $\mu(N)\mbox{ on }\overline{\omega}_{\delta}$
for some positive constant $\mu(N)$ satisfying $\lim_{N\rightarrow+\infty}\mu(N)=+\infty$. We then have
(4.14) $G^{ij}\Psi_{ij}=$ $G^{ij}v_{ij}-tG^{ij}d_{ij}+NG^{ij}d_{i}d_{j}+NdG^{ij}d_{ij}$
$\geqslant$ $(k-l)\psi^{1-1/{(k-l)}}\mu(N)+\eta_{0}\sum_{i}G^{ii}+(Nd-t)G^{ij}d_{ij}$
$\geqslant$ $2\mu_{1}(N)+(\eta_{0}-CN\delta-Ct)\sum_{i}G^{ii}$
on $\overline{\omega}_{\delta}$, where $\mu_{1}(N):=(k-l)(\inf\psi)^{1-1/(k-l)}\mu(N)/2$. Define
(4.15) $\tilde{W}:=1-\exp\{-bW\}.$
By (4.8), we can choose the constant $b$ large enough so that
(4.16) $L\tilde{W}=$ $G^{ij}\big{(}-e^{-bW}b^{2}W_{i}W_{j}+be^{-bW}W_{ij}\big{)}$
$\leqslant$ $be^{-bW}\left[C\left(1+|DW|+\sum_{i}G^{ii}\right)+(C-b)G^{ij}W_{i}W_{j}\right]$
$\leqslant$ $C(1+|D\tilde{W}|+\sum_{i}G^{ii})+(C-b)G^{ij}W_{i}W_{j}be^{-bW}$
$\leqslant$ $C(1+|D\tilde{W}|+\sum_{i}G^{ii}).$
We consider the function
$\Phi:=R\Psi-\tilde{W},$
where $R$ is a large undetermined positive constant. We shall prove
(4.17) Φ0 on $\Phi\leqslant 0\mbox{ on }\overline{\omega}_{\delta}$
by choosing suitable positive constants $\delta$, $t$, $N$ and $R$.
We first consider the case that the maximum of $\Phi$ is achieved at an interior point $x_{0}\in\omega_{\delta}$. It follows that at $x_{0}$,
$|D\tilde{W}|=R|D\Psi|$
and if $N$ is sufficiently large and $\delta<\sqrt{\mu_{1}(N)}/2CN$,
$|D\Psi|=|Dv-tDd+NdDd|\leqslant C(1+t)+C\delta N\leqslant\mu_{1}(N)^{1/2}\mbox{in }\omega_{\delta}.$
Therefore, by (4.14), provided $\delta$ and $t$ sufficiently small such that $CN\delta+Ct<\eta_{0}/2$, we have
(4.18) $L\Psi\geqslant\mu_{1}(N)+\mu_{1}(N)^{1/2}|D\Psi|+\frac{\eta_{0}}{2}\sum_{i}G^{ii}.$
By (4.16) and (4.18) we obtain, at $x_{0}$,
$0\geqslant L\Phi\geqslant$ $R\mu_{1}(N)+R\mu_{1}(N)^{1/2}|D\Psi|+\frac{R\eta_{0}}{2}\sum G^{ii}$
$-C\left(1+|D\tilde{W}|+\sum G^{ii}\right)$
$\geqslant$ $R\mu_{1}(N)-C+R(\mu_{1}(N)^{1/2}-C)|D\Psi|$
$+\left(\frac{R\eta_{0}}{2}-C\right)\sum G^{ii}>0$
provided $N$ and $R$ are chosen sufficiently large which is a contradiction. Thus, the maximum of $\Phi$ is achieved at the boundary $\partial\omega_{\delta}$. We may further assume $\delta<2t/N$ so that
(4.19) $-td+\frac{N}{2}d^{2}\leqslant 0\mbox{ on }\overline{\omega}_{\delta}.$
By (4.11) and (4.19), we can conclude that $\Phi\leqslant 0$ on $\partial\omega_{\delta}$ by choosing $R$ larger and then (4.17) is proved.
Since $(R\Psi-\tilde{W})(0)=0$, we have $(R\Psi-\tilde{W})_{n}(0)\leqslant 0$. Therefore, we get
$u_{n\alpha}(0)\geqslant-C.$
The above arguments also hold for
$W=-\nabla^{\prime}_{\alpha}u-\frac{1}{2}\sum_{1\leqslant\beta\leqslant n-1}u_{\beta}^{2}.$
Hence, we obtain (4.4).
Part 3. the double normal derivative estimate on $\partial\Omega$, that is
(4.20) $|u_{nn}(0)|\leqslant C.$
At the origin, we may assume $\{u_{\alpha\beta}\}_{1\leqslant\alpha,\beta\leqslant n-1}$ is diagonal by a rotation of axes if necessary. Note that
$g^{ij}=\delta_{ij}+\frac{|Du|^{2}}{w^{2}}\delta_{in}\delta_{jn}$
at the origin, we have
$a_{ij}=\begin{pmatrix}\frac{u_{11}}{\omega}&0&\cdots&0&\frac{u_{1n}}{\omega}\\ 0&\frac{u_{22}}{\omega}&\cdots&0&\frac{u_{2n}}{\omega}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\frac{u_{n-1,n-1}}{\omega}&\frac{u_{n-1,n}}{\omega}\\ \frac{u_{n1}}{\omega}&\frac{u_{n2}}{\omega}&\cdots&\frac{u_{n,n-1}}{\omega}&\frac{u_{nn}}{\omega^{3}}\end{pmatrix}$
By Lemma 1.2 of [3] and the estimates of $u_{\alpha\beta}(0)$, $u_{\alpha n}(0)$, there exists a constant $R_{1}>0$ sufficiently large such that if $u_{nn}>R_{1}$, then
$\begin{cases}\kappa_{\alpha}=\dfrac{u_{\alpha\alpha}}{\omega}+o(1),&\alpha=1,\cdots,n-1,\\[10.0pt] \kappa_{n}=\dfrac{u_{nn}}{\omega^{3}}+O(1),\end{cases}$
where $\kappa_{1},\ldots,\kappa_{n}$ are the eigenvalues of $\{a_{ij}\}$.
Since k ¡ n, by choosing $R_{1}$ large enough, we have at the origin,
$\psi=\Big{(}\frac{\sigma_{k}}{\sigma_{l}}\Big{)}(\lambda(\eta))=\frac{\big{(}C_{n}^{k}-(n-k+1)\big{)}u_{nn}^{k}/(w^{3k})+O(u_{nn}^{k-1})}{(C_{n}^{l}-(n-l+1))u_{nn}^{l}/(w^{3l})+O(u_{nn}^{l-1})}.$
It then implies the uniform upper bound of $|u_{nn}(0)|$. Combining (4.3) and (4.4), we can obtain the uniform upper bound of $|u_{ij}(0)|$. Then (4.20) follows immediately and Theorem 4.1 is proved.
It remains to prove Lemma 4.3.
Proof of Lemma 4.3.
By differentiating the equation (4.5), we have
(4.21) $LW+G^{s}W_{s}\leqslant$ $C+C\sum_{i}G^{ii}+C\sum_{s}G^{s}+2G^{ij}u_{ni}\rho_{\alpha j}$
$-\sum_{\beta\leqslant n-1}G^{ij}u_{\beta i}u_{\beta j}.$
By
$a_{ij}=\frac{1}{w}\gamma^{ik}u_{kl}\gamma^{lj},$
we can obtain
$G^{ij}=\frac{\partial G}{\partial u_{ij}}=F^{st}\frac{\partial a_{st}}{\partial u_{ij}}=\frac{1}{w}\sum_{s,t}F^{st}\gamma^{is}\gamma^{tj}$
and
$u_{ij}=w\sum_{s,t}\gamma_{is}a_{st}\gamma_{tj}.$
It follows that
$\sum_{\beta\leqslant n-1}G^{ij}u_{\beta i}u_{\beta j}=w\sum_{\beta\leqslant n-1}\sum_{s,t}F^{ij}\gamma_{\beta s}\gamma_{\beta t}a_{si}a_{tj}$
and
$\frac{1}{w}\sum_{i}F^{ii}\leqslant\sum_{i}G^{ii}\leqslant\frac{1}{w^{3}}\sum_{i}F^{ii}.$
By [13], we can find an orthogonal matrix $B=(b_{ij})$ which can diagonalize $(a_{ij})$ and $(F^{ij})$ at the same time:
$F^{ij}=\sum_{s}b_{is}f_{s}b_{js}\mbox{ and }a_{ij}=\sum_{s}b_{is}\kappa_{s}b_{js}.$
Therefore, we arrive
$\sum_{\beta\leqslant n-1}G^{ij}u_{\beta i}u_{\beta j}=$ $w\sum_{\beta\leqslant n-1}F^{ij}\gamma_{\beta s}\gamma_{\beta t}a_{si}a_{tj}$
$=$ $w\sum_{\beta\leqslant n-1}\sum_{i}\left(\sum_{s}\gamma_{\beta s}b_{si}\right)^{2}f_{i}\kappa_{i}^{2}.$
Let the matrix $\eta=(\eta_{ij})=(\sum_{s}\gamma_{is}b_{sj})$. We can easy to know $\eta\cdot\eta^{T}=g$ and $|\det(\eta)|=\sqrt{1-|Du|^{2}}$. Hence, we obtain
(4.22) $\sum_{\beta\leqslant n-1}G^{ij}u_{\beta i}u_{\beta j}=w\sum_{\beta\leqslant n-1}\sum_{i}\eta_{\beta i}^{2}f_{i}\kappa_{i}^{2}.$
We have
(4.23) $G^{ij}u_{ni}\rho_{\alpha j}=\sum_{i,t}f_{i}\kappa_{i}b_{si}\gamma^{js}b_{ti}\gamma_{nt}\rho_{\alpha j}\leqslant C\sum_{i}f_{i}|\kappa_{i}|.$
For any indices $j,t$, we have
$F^{ij}a_{it}=\sum_{i,s,p}b_{is}f_{s}b_{js}b_{ip}\kappa_{p}b_{tp}=\sum_{i}f_{i}\kappa_{i}b_{ji}b_{ti}.$
Thus, by (4.7), we find
$\mid\sum_{s}G^{s}\mid=$ $\bigg{|}\sum_{s}\biggl{\{}\frac{u_{s}}{w^{2}}\sum_{i}f_{i}\kappa_{i}+\frac{2}{w(1+w)}\sum_{t,j}F^{ij}a_{it}\big{(}wu_{t}\gamma^{sj}+u_{j}\gamma^{ts}\big{)}\biggr{\}}\bigg{|}$
$\leqslant$ $C\sum_{i}f_{i}|\kappa_{i}|.$
Combining (4.21)-(4), we obtain
(4.24) $LW+G^{s}W_{s}\leqslant C\left(1+\sum_{i}G^{ii}+\sum_{i}f_{i}|\kappa_{i}|\right)-w\sum_{\beta\leqslant n-1}\sum_{i}\eta_{\beta i}^{2}f_{i}\kappa_{i}^{2}.$
Now we consider the term $G^{s}W_{s}$. We have, by (4.7) and the definition of the matrix $(b_{ij})$,
(4.25) $-G^{s}W_{s}=$ $-\frac{u_{s}}{w^{2}}\sum_{i}f_{i}\kappa_{i}W_{s}-\frac{2}{w(1+w)}\sum_{t,j}F^{ij}a_{it}\big{(}wu_{t}\gamma^{sj}+u_{j}\gamma^{ts}\big{)}W_{s}$
$=$ $-\frac{1}{w}\sum_{s}\left((k-l)\psi u_{s}\frac{1}{w}+2\sum_{t,i}f_{i}\kappa_{i}(b_{ti}u_{t})\gamma^{sl}b_{li}\right)W_{s}$
$\leqslant$ $C|DW|-\frac{2}{w}\sum_{t,i}f_{i}\kappa_{i}(b_{ti}u_{t})\gamma^{sl}b_{li}W_{s}.$
We further discuss in two situations: (a) $\sum_{\beta\leqslant n-1}\eta_{\beta i}^{2}\geqslant\epsilon$ for all $i$; and (b) $\sum_{\beta\leqslant n-1}\eta_{\beta r}^{2}<\epsilon$ for some index $1\leqslant r\leqslant n$, where $\epsilon$ will be chosen later.
Case (a). By (4.22), we have
$\sum_{\beta\leqslant n-1}G^{ij}u_{\beta i}u_{\beta j}\geqslant\epsilon w\sum_{i}f_{i}\kappa_{i}^{2}.$
By Cauchy-Schwarz inequality, for any $\epsilon_{0}>0$, we have
(4.26) $\frac{2}{w}\kappa_{i}(b_{ti}u_{t})\gamma^{sl}b_{li}W_{s}\geqslant-\frac{\epsilon_{0}}{2}\kappa_{i}^{2}-\frac{C}{\epsilon_{0}}(\gamma^{sl}b_{li}W_{s})^{2}.$
Then
$-G^{s}W_{s}\leqslant C|DW|+\frac{\epsilon_{0}}{2}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon_{0}}G^{ij}W_{i}W_{j}.$
Obviously, for any $\epsilon_{1}>0$,
$\sum_{i}f_{i}|\kappa_{i}|\leqslant\frac{1}{2\epsilon_{1}}\sum_{i}f_{i}+\frac{\epsilon_{1}}{2}\sum_{i}f_{i}\kappa_{i}^{2}\leqslant C\left(\frac{1}{\epsilon_{1}}\sum_{i}G^{ii}+\epsilon_{1}\sum_{i}f_{i}\kappa_{i}^{2}\right).$
Combining the previous four inequalities with (4.24), (4.8) follows.
Case (b). By lemma 4.3 of Bayard [2] , for any $i\neq r$,
$\sum_{\beta\leqslant n-1}\eta^{2}_{\beta i}\geqslant c_{1}$
for some positive constant $c_{1}$ depending on $\|u\|_{C^{1}(\overline{\Omega})}$ and $n$.
Then, in view of (4.22),
(4.27) $\sum_{\beta\leqslant n-1}G^{ij}u_{\beta i}u_{\beta j}\geqslant wc_{1}\sum_{i\neq r}f_{i}\kappa_{i}^{2}.$
If $\kappa_{r}\leqslant 0$, by Lemma 2.7 of [12], we have
$\sum_{i\neq r}f_{i}\kappa_{i}^{2}\geqslant\frac{1}{n+1}\sum_{i=1}^{n}f_{i}\kappa_{i}^{2}.$
Thus (4.8) follows using a similar argument as the Case (a).
So next, we may assume $\kappa_{r}>0$. Without loss of generality, assume that $\kappa_{1}\geqslant\cdots\geqslant\kappa_{n}$. Let $\lambda=(\lambda_{1},\cdots,\lambda_{n})$ be the eigenvalues of $\eta[M_{u}]$, then $\lambda_{i}=\sum_{j\neq i}\kappa_{j}$. It follows that $\lambda_{1}\leqslant\cdots\leqslant\lambda_{n}$. Then we consider two subcases.
Case (b-1). $\kappa_{n}\geqslant-\epsilon_{0}\kappa_{r}$, where $\epsilon_{0}$ is a positive constant to be chosen later.
In order to derive (4.8), the key is to prove the important inequality
(4.28) $f_{r}\kappa_{r}\leqslant C.$
Proof of (4.28): If $\kappa_{n}\geqslant-\epsilon_{0}\kappa_{r}$, then for any $i\neq r$, by choosing $\epsilon_{0}\leqslant\frac{1}{2(n-2)}$ we derive
(4.29) $\lambda_{i}=\sum_{j\neq i}\kappa_{j}\geqslant[1-(n-2)\epsilon_{0}]\kappa_{r}\geqslant\frac{1}{2}\kappa_{r}>0.$
When $r\neq 1$, it is obvious that
$\lambda_{n}\geqslant\cdots\geqslant\lambda_{1}\geqslant[1-(n-2)\epsilon_{0}]\kappa_{r}\geqslant\frac{1}{2}\kappa_{r}>0.$
Thus
(4.30) $f_{r}\kappa_{r}$ $=\sum_{p\neq r}\frac{\sigma_{k-1;p}(\eta)\sigma_{l}(\eta)-\sigma_{k}(\eta)\sigma_{l-1;p}(\eta)}{\sigma_{l}(\eta)^{2}}\kappa_{r}$
$\leqslant(n-k+1)\frac{\sigma_{k-1}(\eta)}{\sigma_{l}(\eta)}\kappa_{r}$
$\leqslant\frac{(n-k+1)C_{n}^{k-1}}{\sigma_{l}(\eta)}\lambda_{n}\cdots\lambda_{n-k+2}\kappa_{r}$
$\leqslant\frac{2(n-k+1)C_{n}^{k-1}}{\sigma_{l}(\eta)}\lambda_{n}\cdots\lambda_{n-k+2}\lambda_{n-k+1}$
$\leqslant C\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}$
$=C\psi.$
When $r=1$, i.e., $\kappa_{n}\geqslant-\epsilon_{0}\kappa_{1}$. By (4.29),
$\lambda_{i}\geqslant[1-(n-2)\epsilon_{0}]\kappa_{1}>0,\quad\forall~{}i\geqslant 2.$
Next we consider two cases: Case $\lambda_{1}\geqslant 0$ and Case $\lambda_{1}<0$.
If $\lambda_{1}\geqslant 0$, similar to (4.30), it is easy to derive $f_{r}\kappa_{r}\leqslant C\psi$.
If $\lambda_{1}<0$, then
$0>\lambda_{1}=\sum_{i\geqslant 2}\kappa_{i}\geqslant(n-1)\kappa_{n}\geqslant-(n-1)\epsilon_{0}\kappa_{1}.$
By $\lambda_{1}<0$, we get $\sigma_{k-1}(\eta)\leqslant\sigma_{k-1;1}(\eta)$. Thus
(4.31) $f_{1}\kappa_{1}$ $\leqslant(n-k+1)\frac{\sigma_{k-1}(\eta)}{\sigma_{l}(\eta)}\kappa_{1}$
$\leqslant(n-k+1)\frac{\sigma_{k-1;1}(\eta)}{\sigma_{l}(\eta)}\kappa_{1}$
$\leqslant\frac{(n-k+1)C_{n-1}^{k-1}}{[1-(n-2)\epsilon_{0}]\sigma_{l}(\eta)}\lambda_{n}\cdots\lambda_{n-k+2}\lambda_{n-k+1}.$
Note that
(4.32) $\sigma_{k}(\eta)$ $=\sigma_{k;1}(\eta)+\lambda_{1}\sigma_{k-1;1}(\eta)$
$\geqslant\lambda_{n}\cdots\lambda_{n-k+1}+\lambda_{1}\cdot C_{n-1}^{k-1}\lambda_{n}\cdots\lambda_{n-k+2}$
$=(\lambda_{n-k+1}+C_{n-1}^{k-1}\lambda_{1})\lambda_{n}\cdots\lambda_{n-k+2}.$
and
$\lambda_{n-k+1}\geqslant[1-(n-2)\epsilon_{0}]\kappa_{1}\geqslant-\frac{1-(n-2)\epsilon_{0}}{(n-1)\epsilon_{0}}\lambda_{1},$
since $n-k+1\geqslant 2$. Therefore by choosing $\epsilon_{0}\leqslant\frac{1}{n-2+2(n-1)C_{n-1}^{k-1}}$ sufficiently small, we get
(4.33) $\frac{1}{2}\lambda_{n-k+1}+C_{n-1}^{k-1}\lambda_{1}\geqslant-\lambda_{1}\left(\frac{1-(n-2)\epsilon_{0}}{2(n-1)\epsilon_{0}}-C_{n-1}^{k-1}\right)\geqslant 0.$
Which implies $\sigma_{k}(\eta)\geqslant\frac{1}{2}\lambda_{n}\cdots\lambda_{n-k+1}$. Combining with (4.31)-(4.33), (4.28) still holds.
By (4.25) and (4.26) and the Cauchy-Schwarz inequality, we have, for any $\epsilon>0$,
(4.34) $-G^{s}W_{s}\leqslant$ $C|DW|-\frac{2}{w}\sum_{i\neq r}f_{i}\kappa_{i}(b_{ti}u_{t})\gamma^{sl}b_{li}W_{s}$
$\leqslant$ $C|DW|+C\sum_{i\neq r}f_{i}|\kappa_{i}||\gamma^{sl}b_{li}W_{s}|$
$\leqslant$ $C|DW|+\epsilon\sum_{i\neq r}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon}\sum_{i=1}^{n}f_{i}\gamma^{sl}b_{li}W_{s}\gamma^{tk}b_{ki}W_{t}$
$\leqslant$ $C|DW|+\epsilon\sum_{i\neq r}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon}G^{ij}W_{i}W_{j}.$
By using (4.27) and fixing $\epsilon$ sufficiently small, we can prove (4.8).
Case (b-2). $\kappa_{n}<-\epsilon_{0}\kappa_{r}$.
Since $\kappa_{r}>0$, it is obvious that $r\neq n$. In this subcase, we get
$\gamma^{sj}b_{jr}W_{s}$ $=\gamma^{sj}b_{jr}(u_{\alpha s}+\rho_{\alpha s}u_{n}+\rho_{\alpha}u_{ns}-\sum_{\beta\leqslant n-1}u_{\beta}u_{\beta s})$
$=\omega\kappa_{r}(\eta_{\alpha r}+\rho_{\alpha}\eta_{nr}-\sum_{\beta\leqslant n-1}u_{\beta}\eta_{\beta r})+\gamma^{sj}b_{jr}\rho_{\alpha s}u_{n}.$
It follows that
(4.35) $|\gamma^{sj}b_{jr}W_{s}|\leqslant C\omega\kappa_{r}(\epsilon+|\rho_{\alpha}|)+C.$
It is obvious that $f_{r}\kappa_{r}=(k-l)\psi-\sum_{i\neq r}f_{i}\kappa_{i}$, thus by $\kappa_{n}^{2}>\epsilon_{0}^{2}\kappa_{r}^{2}$,
$\frac{2}{\omega}f_{r}\kappa_{r}|(\sum_{t}b_{tr}u_{t})\gamma^{sj}b_{jr}W_{s}|=$ $\frac{2}{\omega}((k-l)\psi-\sum_{i\neq r}f_{i}\kappa_{i})\left|\left(\sum_{t}b_{tr}u_{t}\right)\gamma^{sj}b_{jr}W_{s}\right|$
$\leqslant$ $C|DW|+C(\epsilon+|\rho_{\alpha}|)\sum_{i\neq r}f_{i}|\kappa_{i}|\kappa_{r}+C\sum_{i\neq r}f_{i}|\kappa_{i}|$
$\leqslant$ $C|DW|+C\epsilon_{0}^{-1}(\epsilon+|\rho_{\alpha}|)\sum_{i\neq r}f_{i}\kappa_{i}^{2}$
$+C\epsilon_{0}(\epsilon+|\rho_{\alpha}|)\sum_{i\neq r}f_{i}\kappa_{r}^{2}+C\epsilon_{1}\sum_{i\neq r}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon_{1}}\sum_{i\neq r}f_{i}$
$\leqslant$ $C\psi|DW|+(C\epsilon_{0}^{-1}(\epsilon+|\rho_{\alpha}|)+\epsilon_{1})\sum_{i\neq r}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon_{1}}\sum_{i}G^{ii}$
for any $\epsilon_{1}>0$. Here the last inequality comes from $\sum_{i\neq r}f_{i}\kappa_{r}^{2}\leqslant\frac{n-1}{\epsilon_{0}^{2}}f_{n}\kappa_{n}^{2}\leqslant\frac{n-1}{\epsilon_{0}^{2}}\sum_{i\neq r}f_{i}\kappa_{i}^{2}$ with $r\neq n$.
We can choose sufficiently small $\delta$, $\epsilon$ and $\epsilon_{1}$ satisfying
$\left(\epsilon_{0}^{-1}C(\epsilon+|\rho_{\alpha}|)+\epsilon_{1}\right)<\frac{mc_{1}}{4},$
where $m=1-\theta^{2}$, $\theta$ is defined in spacelike. Therefore, as in Case (b-1), (4.8) follows by (4.27). ∎
The last step is to estimate the second derivatives in the interior.
We need the following lemma of Urbas[28]. As usual, unless otherwise specified, we assume that the same indicator represents the summation.
Lemma 4.4.
The second fundamental form $h_{ab}$ satisfies
(4.36) $F^{ij}\nabla_{i}\nabla_{j}h_{ab}=$ $-F^{ij,kl}\nabla_{a}h_{ij}\nabla_{b}h_{kl}-F^{ij}h_{ij}h_{ak}h_{bk}$
$+F^{ij}h_{ik}h_{jk}h_{ab}+\nabla_{a}\nabla_{b}\psi$
where
$F^{ij}=\frac{\partial f(\lambda(h))}{h_{ij}}.$
Suppose the maximum of $h_{\xi\xi}$ is achieved at $X_{0}=(x_{0},u(x_{0}))\in M_{u}$ and $\xi_{0}\in T_{X_{0}}M_{u}$. We take point $X_{0}$ as the origin and establish a local orthonormal frame $\{e_{1},e_{2},\cdots,e_{n}\}$ such that,
$\xi_{0}=e_{1},\ \nabla_{e_{i}}e_{j}=0\mbox{ at }X_{0}.$
We may also assume that $\{h_{ij}\}$ is diagonal at $X_{0}$ and furthermore,
$h_{11}\geqslant\cdots\geqslant h_{nn}$
by rotating the coordinate system if necessary.
Using (4.36) and differentiating the equation (1.1) twice, we have
(4.37) $0\geqslant\tilde{F}^{ii}\nabla_{ii}h_{11}=-\tilde{F}^{ij,kl}\nabla_{1}h_{ij}\nabla_{1}h_{kl}-\tilde{\psi}h_{11}^{2}+h_{11}\tilde{F}^{ii}h_{ii}^{2}+\nabla_{11}(\tilde{\psi}).$
where $\tilde{\psi}=\psi^{1/(k-l)}$ and
$\tilde{F}^{ij}=\frac{\partial(\frac{\sigma_{k}}{\sigma_{l}})^{1/(k-l)}(\lambda(h))}{\partial h_{ij}},\;\;\;\tilde{F}^{ij,kl}=\frac{\partial^{2}(\frac{\sigma_{k}}{\sigma_{l}})^{1/(k-l)}(\lambda(h))}{\partial h_{ij}\partial h_{kl}}.$
Thus, it follows from (2.6) and the concavity of $(\frac{\sigma_{k}}{\sigma_{l}})^{1/(k-l)}(\lambda(h))$ that
$0\geqslant c(n,k,l)h_{11}^{3}-\tilde{\psi}h_{11}^{2}+\nabla_{11}(\tilde{\psi})\geqslant c(n,k,l)h_{11}^{3}-Ch_{11}^{2}$
provided $h_{{11}}$ large enough. Then a bound for $h_{11}$ is derived from the above inequality.
Combined with the boundary estimates of the second-order derivatives (4.1), we have
(4.38) $\max_{\overline{\Omega}}|D^{2}u|\leqslant C.$
Proof of theorem 1.3.
Combined (3.2), (3.1), (4.38), we can obtain
(4.39) $||u||_{C^{2}(\overline{\Omega})}\leqslant C.$
Applying the Evans-Krylov theory, we can obtain the $C^{2,\alpha}$ estimates for $u$. Using the continuity method and the maximum principle, we can finally establish the existence and uniqueness of solutions to (1.3). ∎

### v2

@@
 Abstract.
-In this paper, we consider the Dirichlet problem for a class of prescribed Hessian quotient type curvature equations in Minkowski space.
-By establishing the a priori $C^{2}$ estimates, we prove the existence of smooth spacelike hypersurfaces with a class of prescribed Hessian quotient type curvature and homogeneous boundary data in the non-degenerate case.
+In this paper, we consider the Dirichlet problem for a class of prescribed Hessian quotient type curvature equations with homogeneous boundary data in Minkowski space.
+By establishing the a priori $C^{2}$ estimates, we obtain the existence result in the non-degenerate case.
 Keywords: Minkowski space; Hessian quotient type curvature equations; a priori $C^{2}$ estimates.
 1. Introduction
-In this paper, we consider the Dirichlet problem for the following non-degenerate Hessian quotient type curvature equations in Minkowski space $\mathbb{R}^{n,1}$:
-(1.1) $\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))=\psi(X).$
-When $k=n$ and $l=0$, this problem was studied by H. Jiao and the first author in [15] for general boundary data.
+In Minkowski space $\mathbb{R}^{n,1}$, a function $u$ defined on a bounded domain $\Omega\subset\mathbb{R}^{n}$ is called spacelike if
+$\sup_{\overline{\Omega}}|Du|<1.$
+In this paper, we are interested in the prescribed curvature problems for graph $M_{u}$ of a spacelike function $u$, which is
+$M_{u}:=\{(x,u(x)):x\in\Omega\subset\mathbb{R}^{n}\}.$
+We shall consider the Dirichlet problem for the following non-degenerate Hessian quotient type curvature equations:
+(1.1) $\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))=\psi(X)\;\mbox{in}\;\Omega$
+with homogeneous boundary data.
+Here $X=(x,u(x))$ is the position vector, non-degenerate means $\psi(X)>0$, $\sigma_{k}$ is the $k$-th elementary symmetric function, $\lambda(\eta[M_{u}])=(\lambda_{1},\cdots,\lambda_{n})$ denotes the eigenvalues of $\eta$ with respect to $g$ on $M_{u}$.
+$\eta$ is the $(0,2)$-tensor field defined on $M$ by
+$\eta=Hg-h,$
+where $H(X)$ is the mean curvature of $M_{u}$ at $X\in M_{u}$, $g=\{g_{ij}\}$ is the induced Riemannian metric and $h=\{h_{ij}\}$ is the second fundamental form.
+When $k=n$ and $l=0$, this problem was studied by Jiao and the first author in [16] for general boundary data.
 In Euclidean space, Chen-Tu-Xiang [10] derived the existence and uniqueness results for the Dirichlet problem of
 $\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))=\psi(X,\nu),\;\;k<n$
-with homogeneous boundary data in degenerate form. The current work aims to extend the results of [9] from Euclidean space to Minkowski space $\mathbb{R}^{n,1}$, whose induced metric is given by
+with homogeneous boundary data in the degenerate case. The current work aims to extend the results of [10] from Euclidean space to Minkowski space $\mathbb{R}^{n,1}$, whose induced metric is given by
 $ds^{2}=dx^{2}_{1}+\cdots+dx^{2}_{n}-dx^{2}_{n+1}.$
-Due to the lack of gradient estimates, we restrict our analysis to the non-degenerate case $\psi=\psi(X)>0$, see Remark 3.2.
-In contrast to Euclidean space, when differentiating equations twice in the Minkowski space, uncontrollable terms including the square of curvatures will appear.
-This is the reason why the Dirichlet problem for $k$-curvature equations
+In contrast to the Euclidean case, when differentiating equations twice in the Minkowski space, uncontrollable terms including the square of curvatures will appear.
+Hence the Dirichlet problem for $k$-curvature equations
 (1.2) $\sigma_{k}(\kappa[M_{u}])=\psi$
-remains unsolved for $33\leqslant k\leqslant n-$ 3, although it has been solved a long time ago for corresponding equations in Euclidean space, where
+remains unsolved for $3\leqslant k\leqslant n-3$, although it has been solved a long time ago for its Euclidean analogue, where
 $\sigma_{k}(\kappa[M_{u}])=\sum_{1\leqslant i_{1}<\cdots<i_{k}\leqslant n}\kappa_{i_{1}}\cdots\kappa_{i_{k}},$
-where $\kappa[M_{u}]=(\kappa_{1},\cdots,\kappa_{n})$ denote the principal curvatures of $M_{u}$.
-The reader is referred to [4, 16, 17, 22, 18, 6, 7, 30] for the study of the Dirichlet problem for (1.2) and the quotient type curvature equation in Euclidean space.
-In this paper, we are interested in the graph $M_{u}$ of a spacelike function $u$ defined on a bounded domain $\Omega\subset\mathbb{R}^{n}$, which is
-$M_{u}:=\{(x,u(x)):x\in\Omega\subset\mathbb{R}^{n}\}.$
-For the function $v(x)\in C^{2}(\Omega)$, we denote $v_{i}=\frac{\partial v}{\partial x_{i}}$, $v_{ij}=\frac{\partial^{2}v}{\partial x_{i}\partial x_{j}}$, and $Dv=(v_{1},\cdots,v_{n})$ and $D^{2}v=(v_{ij})$ represent the ordinary gradient and Hessian matrix.
-In Minkowski space, we call a function $u$ spacelike if
-$\sup_{\overline{\Omega}}|Du|<1.$
+$\kappa[M_{u}]=(\kappa_{1},\cdots,\kappa_{n})$ denotes the principal curvatures of $M_{u}$.
+The reader is referred to [5, 7, 8, 15, 17, 18, 19, 23, 25, 33] for the study of the Dirichlet problem for (1.2) and the quotient type curvature equation in Euclidean space.
+For the function $v(x)\in C^{2}(\Omega)$, we denote $v_{i}=\frac{\partial v}{\partial x_{i}}$, $v_{ij}=\frac{\partial^{2}v}{\partial x_{i}\partial x_{j}}$, and use $Dv=(v_{1},\cdots,v_{n})$ and $D^{2}v=\{v_{ij}\}$ to represent the ordinary gradient and Hessian matrix respectively.
 Clearly, for the spacelike function $u$, the Minkowski metric restricted to $M_{u}$ induces a Riemannian metric on $M_{u}$:
-[…]
+$g_{ij}=\delta_{ij}-u_{i}u_{j},\ 1\leqslant i,j\leqslant n.$
 The inverse of the metric is
-[…]
+$g^{ij}=\delta_{ij}+\frac{u_{i}u_{j}}{1-|Du|^{2}},\ 1\leqslant i,j\leqslant n.$
 The second fundamental form $M_{u}$ with respect to its upward unit normal vector field
@@
 $h_{ij}=\frac{u_{ij}}{\sqrt{1-|Du|^{2}}}.$
-Define the $(0,2)$-tensor field $\eta$ on $M$ by
-$\eta=Hg-h,$
-where $H(X)$ is the mean curvature of $M_{u}$ at $X=(x,u(x))\in M_{u}$, $g=\{g_{ij}\}$, $h=\{h_{ij}\}$.
-$\lambda(\eta[M_{u}])=(\lambda_{1},\cdots,\lambda_{n})$ denote the eigenvalues of $\eta$ with respect to $g$ on $M_{u}$.
-In this paper, we consider the existence of smooth spacelike graphic hypersurface satisfying the Dirichlet problem
-(1.3) $\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))&=\psi(x,u)&\;\;~{}\mbox{ in }\Omega,\\ u&=0&\;\;~{}\mbox{ on }\partial\Omega,\end{aligned}\right.$
+In this paper, we consider the existence of smooth spacelike graphic hypersurfaces satisfying the Dirichlet problem
+(1.3) $\left\{\begin{aligned} \frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))&=\psi(x,u)&\;\;~\mbox{ in }\Omega,\\ u&=0&\;\;~\mbox{ on }\partial\Omega,\end{aligned}\right$.
 where $\psi\in C^{\infty}(\overline{\Omega}\times\mathbb{R})>0$.
-We first introduce the definition of admissible set of equation (1.1).
+Define the Gårding’s cone
+$\Gamma_{k}:=\{\lambda=(\lambda_{1},\cdots,\lambda_{n})\in\mathbb{R}^{n}:\sigma_{m}(\lambda)>0,m=1,\cdots,k\}.$
+We now introduce the definition of admissible set of equation (1.1).
 Definition 1.1.
-A $C^{2}$ regular spacelike hypersurface $M_{u}$ is called $(\eta,k)$-convex if the principal curvatures $\kappa=(\kappa_{1},\ldots,\kappa_{n})\in\widetilde{\Gamma}_{k}$ at each $X\in M_{u}$, where $\widetilde{\Gamma}_{k}$ is the symmetric cone defined by
+A $C^{2}$ regular spacelike hypersurface $M$ is called $(\eta,k)$-convex if the principal curvatures $\kappa=(\kappa_{1},\ldots,\kappa_{n})\in\widetilde{\Gamma}_{k}$ at each $X\in M$, where $\widetilde{\Gamma}_{k}$ is the symmetric cone defined by
 (1.4) $\widetilde{\Gamma}_{k}:=\{\kappa=(\kappa_{1},\ldots,\kappa_{n})\in\mathbb{R}^{n}:\lambda(\eta)=(\lambda_{1},\cdots,\lambda_{n})\in\Gamma_{k},\lambda_{i}(\eta)=\sum_{j\neq i}\kappa_{j}\}.$
-$\Gamma_{k}$ is defined by
-$\Gamma_{k}:=\{\lambda=(\lambda_{1},\cdots,\lambda_{n})\in\mathbb{R}^{n}:\sigma_{m}(\lambda)>0,m=1,\cdots,k\}.$
 In addition, we call a $C^{2}$ spacelike function $u:\Omega\rightarrow\mathbb{R}$ admissible if its graph $M_{u}$ is $(\eta,k)$-convex.
-We also need some geometric conditions on $\Omega$ similar to those in [17] when consider the Dirichlet boundary value problem.
-Definition 1.2.
-If $k<n$, a bounded domain $\Omega\subset\mathbb{R}^{n}$ is called $(\eta,k)$-convex, if there exists a positive constant $K$ such that for each $x\in\partial\Omega$,
-$(\kappa^{b}_{1}(x),\ldots,\kappa^{b}_{n-1}(x),K)\in\widetilde{\Gamma}_{k},$
-where $\kappa^{b}_{1}(x),\ldots,\kappa^{b}_{n-1}(x)$ are the principal curvatures of $\partial\Omega$ at $x$.
-Jiao and Sun [20] considered the following degenerate curvature equation
+Jiao-Sun [21] considered the following degenerate curvature equation
 $\det(\lambda(\eta[M_{u}]))=\psi(X,\nu)$
@@
 They derived the existence of $C^{1,1}$ regular graphical $(\eta,n)$-convex hypersurfaces.
-Motivated by [20], Chen-Tu-Xiang [9] extended this work to a degenerate Hessian quotient type curvature equation (1.1) with $0\leqslant l<k<n$, they proved the existence and uniqueness of $C^{1,1}$ regular graphical hypersurfaces in Euclidean space under the condition $\psi^{\frac{1}{k-1}}(x,u(x),\nu(x))\in C^{1,1}(\Omega\times\mathbb{R}\times S^{n})$.
+Motivated by [21], Chen-Tu-Xiang [10] extended this work to a degenerate Hessian quotient type curvature equation (1.1) with $0\leqslant l<k<n$ and proved the existence and uniqueness of $C^{1,1}$ regular graphical hypersurfaces in Euclidean space under the condition $\psi^{\frac{1}{k-1}}(x,u(x),\nu(x))\in C^{1,1}(\Omega\times\mathbb{R}\times S^{n})$.
 Building on [10], we generalize their results to Minkowski space and establish the following theorem:
@@
 Let $k\geqslant 2,~0\leqslant l<k<n$.
-Assume that bounded domain $\Omega\subset\mathbb{R}^{n}$ is $(\eta,k)$-convex with smooth boundary $\partial\Omega$.
+Assume that $\Omega\subset\mathbb{R}^{n}$ is bounded domain with smooth boundary $\partial\Omega$.
 Suppose $\psi=\psi(x,z)\in C^{\infty}(\overline{\Omega}\times\mathbb{R})>0$, $\psi_{z}\geqslant 0$.
 In addition, assume that there exists an admissible subsolution $\underline{u}\in C^{2}(\overline{\Omega})$ satisfying $\kappa[M_{\underline{u}}]\in\widetilde{\Gamma}_{k}$ and
-[…]
+(1.5) $\left\{\begin{aligned} &\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{\underline{u}}]))\geqslant\psi^{0}(x,\underline{u})\geqslant\psi(x,u),&&in~\Omega,\\ &\underline{u}=0,&&on~\partial\Omega\end{aligned}\right$.
 Then there exists a unique admissible solution $u\in C^{\infty}(\overline{\Omega})$ to (1.3).
+Generally speaking, we need some geometric conditions on $\Omega$ as in [18] to deal with equation (1.1).
+Such convexity-type conditions are typically used to derive subsolutions or construct barriers in second order estimates on the boundary for Hessian equation and curvature equations in the form
+$F(D^{2}u)=f(\lambda(D^{2}u))=\psi\;\mbox{and}\;F(\kappa(M))=\psi.$
+For equations like (1.1), the correspond condition is: there exists a positive constant $K$ such that for each $x\in\partial\Omega$,
+(1.6) $(\kappa_{1}^{\prime}(x),\dots,\kappa_{n-1}^{\prime}(x),K)\in\widetilde{\Gamma}_{k},$
+where $\kappa_{1}^{\prime}(x),\dots,\kappa_{n-1}^{\prime}(x)$ are the principal curvatures of $\partial\Omega$ at $x$. However, we notice that for equation (1.1), when $k<n$,
+$\sigma_{k}(\lambda(\kappa_{1}^{\prime},\cdots\kappa_{n-1}^{\prime},K))=C_{n-1}^{k}K^{k}+\mbox{bounded terms},$
+where $\lambda=(\lambda_{1},\cdots,\lambda_{n})$ is defined as in (1.4).
+It follows that, for any $K$ sufficiently large, $\sigma_{k}(\lambda(\kappa_{1}^{\prime},\cdots\kappa_{n-1}^{\prime},K))$ is always positive and thus,
+$(\kappa_{1}^{\prime}(x),\dots,\kappa_{n-1}^{\prime}(x),K)\in\widetilde{\Gamma}_{q},\;q=1,2,\ldots,n-1.$
+In fact, since the positive $\kappa_{i}$ axes belong to $\tilde{\Gamma}_{k}$ for $k<n$, it is the cone of type 2 (see [4] for the definition) and condition (1.6) holds automatically.
+Therefore, the geometric restrictions on $\Omega$ can be dropped in Theorem 1.2.
 To conclude the introduction, we briefly mention some related studies.
-If $\lambda(\eta[M_{u}])$ is replaced by the principal curvature $\kappa[M_{u}]$ of $M_{u}$ and $l=0$, equation (1.1) becomes the classical $k$-curvature equations (1.2).
-The mean curvature, scalar curvature, and Gauss curvature correspond respectively to $k=1,2,n$ in the formula (1.2).
-The Dirichlet problem for prescribed mean curvature in Minkowski space was first solved by Bartnik and Simon [1], while Delanoè [10] studied the Dirichlet problem of the prescribed Gauss curvature equation (see also [11]).
-Bayard [2] completely solved the solvability of the Dirichlet problem of the prescribed scalar curvature equation in four-dimensional space under the condition of strict convexity, and Urbas [28] generalized the result to the general dimension $n$.
-Ren and Wang [24, 25] solved the case $k=n-1$ and $k=n-2$.
-For $33\leqslant k\leqslant n-$ 3, this is still an open problem.
-The reason is the lack of a priori $C^{2}$ estimates.
-For this problem, Guan-Ren-Wang [14] given a second order estimates for admissible solutions of $k$-curvature equations (1.2) whose graphs are $(k+1)$-convex, and Wang [29] established the estimates for admissible solutions whose graphs are semi-convex.
+If $\lambda(\eta[M_{u}])$ is replaced by the principal curvature $\kappa$ of $M_{u}$ and $l=0$, equation (1.1) becomes the classical $k$-curvature equation (1.2).
+The mean curvature, scalar curvature and Gauss curvature equations correspond respectively to $k=1,2,$ and $n$ in the formula (1.2).
+The Dirichlet problem for the prescribed mean curvature equation in Minkowski space was first solved by Bartnik-Simon [1], while Delanoè [11] studied the Dirichlet problem of the prescribed Gauss curvature equation (see also [12]).
+Bayard [2] solved the Dirichlet problem for the prescribed scalar curvature equation in dimension four under the condition of strict convexity, and Urbas [30] generalized the result to general dimension.
+Ren-Wang [26, 27] solved the case k=n1k=n-1 and k=n2k=n-2.
+For $3\leqslant k\leqslant n-3$, the problem remains open, due to the lack of a priori $C^{2}$ estimates.
+Wang [32] obtained the curvature estimates for (k+1)-convex solutions of $k$-curvature equations (1.2) in Minkowski space, while Wang [31] established the estimates under semi-convex condition.
 For the corresponding Hessian quotient type equations
 $\frac{\sigma_{k}(\lambda(\Delta uI-D^{2}u))}{\sigma_{l}(\lambda(\Delta uI-D^{2}u))}=f(x,u(x),Du(x)),$
-Liu-Mao-Zhao [23] established Pogorelov type estimates of $k$-convex solutions with the vanishing Dirichlet boundary condition in Lorentz-Minkowski space $\mathbb{R}^{n+1}$.
-Chen-Tu-Xiang [8] considered the Dirichlet problem for the above equation on compact Riemannian Manifolds and obtained the existence and uniqueness.
+Liu-Mao-Zhao [24] established Pogorelov type estimates of $k$-convex solutions with homogeneous boundary data in Lorentz-Minkowski space $\mathbb{R}^{n+1}$.
+Chen-Tu-Xiang [9] considered the Dirichlet problem on compact Riemannian manifolds and obtained the existence result.
 The rest of this paper is organized as follows.
-In Section 2, we provided some preliminaries.
-In Section 3, The $C^{1}$ estimates are established.
-Section 4 is devoted to estimates for second order derivatives.
+In Section 2, we provide some preliminaries.
+The $C^{1}$ estimate is estabished in Section 3.
+We deal with the boundary second order derivative estimates in Section 4.
+Finally, the proof of Theorem 1.2 is completed in Section 5.
 2. Preliminaries
@@
 $M_{u}=\{X=(x,u(x))|x\in\mathbb{R}^{n}\}.$
-[…]
+Then $u(x)=-\langle X,\epsilon_{n+1}\rangle$, where $\langle\cdot,\cdot\rangle$ is the Minkowski inner product.
 On the other hand, when we choose local orthonormal frames $\{e_{1},e_{2},\cdots,e_{n}\}$ on $TM_{u}$.
@@
 where $h_{ij}=\langle D_{e_{i}}\nu,e_{j}\rangle$ is the second fundamental form of $M$.
-Define the martix $A=\{a_{ij}\}$, where
-(2.2) $a_{ij}=\frac{1}{w}\gamma^{ik}u_{kl}\gamma^{lj},$
-and $w=\sqrt{1-|Du|^{2}}$, $\gamma^{ik}=\delta_{ik}+\frac{u_{i}u_{k}}{w(1+w)}.$ Note that $\{\gamma^{ij}\}$ is invertible and its inverse is $\{\gamma_{ij}\}=\{\delta_{ij}-\frac{u_{i}u_{j}}{1+w}\}$, which is the square root of $\{g_{ij}\}.$
+Define
+(2.2) $A=\{a_{ij}\}=\Big\{\frac{1}{w}\gamma^{ik}u_{kl}\gamma^{lj}\Big\},$
+where $w=\sqrt{1-|Du|^{2}}$ and $\gamma^{ik}=\delta_{ik}+\frac{u_{i}u_{k}}{w(1+w)}.$ Note that $\{\gamma^{ij}\}$ is invertible and its inverse is $\{\gamma_{ij}\}=\{\delta_{ij}-\frac{u_{i}u_{j}}{1+w}\}$, which is the square root of $\{g_{ij}\}.$
 Clearly, $A=\{a_{ij}\}$ is symmetric and the eigenvalues of $A$ equal the prescribed principal curvatures of $M_{u}$, which means
-$\lambda(\{a_{ij}\})=\lambda(\{g^{ij}\}\{h_{ij}\})=\lambda\big{(}\frac{1}{w}\left(I+\frac{Du\otimes Du}{w^{2}}\right)D^{2}u\big{)}.$
+(2.3) $\kappa(\{a_{ij}\})=\kappa(\{g^{ij}\}\{h_{ij}\})=\kappa\big(\frac{1}{w}\left(I+\frac{Du\otimes Du}{w^{2}}\right)D^{2}u\big).$
 Note that the equation (1.1) can be rewritten by
@@
 Denote $\sigma_{m;i_{1},\cdots,i_{k}}(\kappa)=\sigma_{m}(\kappa)|_{\kappa_{i_{1}}=\cdots=\kappa_{i_{k}}=0}$ for integer $1\leqslant i_{1},\cdots,i_{k}\leqslant n,1\leqslant m\leqslant n$ and $n-k\leqslant m$.
-Then we have:
+Then we have,
 $\frac{\partial\sigma_{k}}{\partial\kappa_{i}}(\kappa)=\sigma_{k-1;i}(\kappa),$
@@
 Proposition 2.1.
-For $\kappa\in\Gamma_{k}$ and $n\geqslant k>l\geqslant 0$, $n\geqslant r>s\geqslant 0$, $k\geqslant r$, $l\geqslant s$, we have
-$\Bigg{[}\frac{{\sigma_{k}(\kappa)}/{C_{n}^{k}}}{{\sigma_{l}(\kappa)}/{C_{n}^{l}}}\Bigg{]}^{\frac{1}{k-l}}\leqslant\Bigg{[}\frac{{\sigma_{r}(\kappa)}/{C_{n}^{r}}}{{\sigma_{s}(\kappa)}/{C_{n}^{s}}}\Bigg{]}^{\frac{1}{r-s}}.$
+For $\lambda\in\Gamma_{k}$ and $n\geqslant k>l\geqslant 0$, $n\geqslant r>s\geqslant 0$, $k\geqslant r$, $l\geqslant s$, we have
+$\Bigg[\frac{{\sigma_{k}(\lambda)}/{C_{n}^{k}}}{{\sigma_{l}(\lambda)}/{C_{n}^{l}}}\Bigg]^{\frac{1}{k-l}}\leqslant\Bigg[\frac{{\sigma_{r}(\lambda)}/{C_{n}^{r}}}{{\sigma_{s}(\lambda)}/{C_{n}^{s}}}\Bigg]^{\frac{1}{r-s}}.$
 Proof.
 See [28]. ∎
-Next, we introduce a new cone $\widetilde{\Gamma}_{k}$ defined by (1.4) and list some important properties, more details can refer to [5].
+Next, we list some important properties cone $\widetilde{\Gamma}_{k}$ defined by (1.4), more details can refer to [6].
 Proposition 2.2.
@@
 (1) $\widetilde{\Gamma}_{k}$ are convex cones and
-(2.4) $\Gamma_{1}=\widetilde{\Gamma}_{1}\supset\widetilde{\Gamma}_{2}\supset\cdots\supset\widetilde{\Gamma}_{n}\supset\Gamma_{2}.$
+(2.5) $\Gamma_{1}=\widetilde{\Gamma}_{1}\supset\widetilde{\Gamma}_{2}\supset\cdots\supset\widetilde{\Gamma}_{n}\supset\Gamma_{2}\supset\Gamma_{3}\cdots\supset\Gamma_{n}.$
 (2) If $\kappa=(\kappa_{1},\cdots,\kappa_{n})\in\widetilde{\Gamma}_{k}$, then $\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]^{\frac{1}{k-l}}(0\leqslant l<k\leqslant n)$ are concave with respect to $\kappa$.
@@
 3. $C^{1}$ estimates
-In this section, we establish the $C^{1}$ estimates for the admissible solution to (1.3).
-Simliar as [15], we prove the admissible solution $u$ is spacelike, which means there exists a constant $10<\theta_{0}<$ 1 such that
+Next, we establish the $C^{1}$ estimates for the admissible solution to (1.3).
+Simliar as [16], we prove the admissible solution $u$ is spacelike, which means there exists a constant $0<\theta_{0}<1$ such that
 (3.1) $\sup_{\overline{\Omega}}|Du|\leqslant 1-\theta_{0}.$
-Since $M_{u}$ is $(\eta,k)$-convex, then $\sigma_{1}(\lambda(\eta[M_{u}]))>0$, hence $H[M_{u}]>0$.
-By proposition (2.1), we can find a spacelike supersolution $\overline{u}\in C^{\infty}(\overline{\Omega})$ of (1.3).
-Then combining with $\psi_{u}\geqslant 0$, the subsolution condition (1.5) and the maximum principle, we can find that
+This global gradient estimates will give an upper bound of $\tilde{w}$, which ensures the uniform ellipticity of $\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta))$ if $u||_{C^{2}(\overline{\Omega})}<C$.
+Since $M_{u}$ is $(\eta,k)$-convex, we have $\sigma_{1}(\lambda(\eta[M_{u}]))>0$, which means $\sigma_{1}(\kappa(M_{u}))=H[M_{u}]>0$.
+Obviously, $\bar{u}=0$ is a supersolution of Dirichlet problem (1.3).
+Combining with $\psi_{u}\geqslant 0$, the subsolution condition (1.5) and the maximum principle, we can find that
+$\underline{u}\leqslant u\leqslant\overline{u}\text{ in }\Omega\text{ and }\overline{u}=u=\underline{u}\text{ on }\partial{\Omega},$
+which implies
+$\frac{\partial\underline{u}}{\partial\gamma}\leqslant\frac{\partial u}{\partial\gamma}\leqslant\frac{\partial\overline{u}}{\partial\gamma},$
+where $\gamma$ is the interior unit normal to $\partial\Omega$. Then we have
 (3.2) $\sup_{\overline{\Omega}}|u|\leqslant\max\{\sup_{\overline{\Omega}}|\overline{u}|,\sup_{\overline{\Omega}}|\underline{u}|\}\leqslant C$
@@
 (3.3) $\sup_{\partial\Omega}|Du|\leqslant\max\{\sup_{\partial\Omega}|D\underline{u}|,\sup_{\partial\Omega}|D\overline{u}|\}\leqslant 1-\theta$
-for some constant $10<\theta<$ 1 since $\underline{u}$ and $\overline{u}$ are both spacelike. Next we establish the global gradient estimates (3.1), that is, we prove that the function
+for some constant $0<\theta<1$ since $\underline{u}$ and $\overline{u}$ are both spacelike.
+Next, we prove an upper bound for
 $\tilde{w}=\frac{1}{\sqrt{1-|Du|^{2}}}=\frac{1}{w}.$
-has an upper bound.
 Theorem 3.1.
 Let $u\in C^{3}(\Omega)\cap C^{1}(\overline{\Omega})$ be an admissible solution of (1.3). Suppose the smooth function $\psi$ satisfies $\psi_{u}\geqslant 0$. Then
-(3.4) $\sup_{\overline{\Omega}}\tilde{w}\leqslant C(1+\sup_{\partial\Omega}\tilde{w}),$
+[…]
 where $C$ depends on $|D\psi|$, the lower bound of $\psi$, $\sup_{\overline{\Omega}}|u|$ and other known data.
@@
 $e_{i}=\gamma^{is}\tilde{\partial}_{s},\ i=1,\ldots,n,$
-where $\gamma^{is}:=\delta_{is}+\frac{u_{i}u_{s}}{w(1+w)}$ and ~$\tilde{\partial}_{s}:=\epsilon_{s}+u_{s}\epsilon_{n+1}$.
-It’s clear that $\{e_{1},e_{2},\cdots,e_{n}\}$ be an orthonormal frame on $TM_{u}$ around $X_{0}=(x_{0},u(x_{0}))$ such that, at $x_{0}$,
-(3.5) $\nabla_{1}u=\frac{|Du|}{w}=|\nabla u|,\nabla_{i}u=u_{i}=0,\text{ for }i\geqslant 2.$
+where $\gamma^{is}:=\delta_{is}+\frac{u_{i}u_{s}}{w(1+w)}$ and $\tilde{\partial}_{s}:=\epsilon_{s}+u_{s}\epsilon_{n+1}$.
+It’s clear that $\{e_{1},e_{2},\cdots,e_{n}\}$ is an orthonormal frame on $TM_{u}$ around $X_{0}=(x_{0},u(x_{0}))$ such that, at $x_{0}$,
+(3.5) $\nabla_{1}u=\frac{|Du|}{w}=|\nabla u|,\nabla_{i}u=\frac{u_{i}}{w}=0,\text{ for }i\geqslant 2.$
 We may further rotate $\epsilon_{2},\ldots,\epsilon_{n}$ such that $\{u_{ij}\}_{i,j\geqslant 2}$ is diagonal at $x_{0}$. At $x_{0}$, we have
@@
 and
-(3.7) $0\geqslant F^{ii}\nabla_{ii}Q=\frac{F^{ii}\nabla_{i}h_{i1}\nabla_{1}u+F^{ii}h_{il}\nabla_{il}u}{\tilde{w}}-\frac{F^{ii}\left(h_{i1}\nabla_{1}u\right)^{2}}{\tilde{w}^{2}}+BF^{ii}\nabla_{ii}u,$
+(3.7) $0\geqslant F^{ij}\nabla_{ij}Q=\frac{F^{ij}\nabla_{j}h_{i1}\nabla_{1}u+F^{ij}h_{il}\nabla_{jl}u}{\tilde{w}}-\frac{F^{ij}(h_{i1}\nabla_{1}u)(h_{j1}\nabla_{1}u)}{\tilde{w}^{2}}+BF^{ij}\nabla_{ij}u,$
 where
@@
 and (3.4) follows as [2]. It follows from (3.3) and (3.4) that (3.1) holds. ∎
-Remark 3.2.
-Although Chen-Tu-Xiang [9] has given the gradient estimate for admissible solutions of the equation
-$\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))=\psi(x,u,\nu)\geqslant 0$
-in Euclidean space, gradient estimates in Minkowski space remain a challenge for degenerate problems due to the metric.
-In addition, no satisfactory approach has been developed to control the term $\nu$, even in non-degenerate cases.
 4. $C^{2}$ estimates
@@
 Theorem 4.1.
-Suppose bounded domain $\Omega\subset\mathbb{R}^{n}$ is $(\eta,k)$-convex with smooth boundary $\partial\Omega$.
+Suppose $\Omega\subset\mathbb{R}^{n}$ is a bounded domain with smooth boundary $\partial\Omega$.
 Let $u\in C^{3}(\overline{\Omega})$ be an admissible solution of (1.3).
-Then there exists a positive constant $C$ depending only on $\theta_{0}$, $\|\psi\|_{C^{1}(\overline{\Omega}\times[-\mu_{0},\mu_{0}])}$, $\|\varphi\|_{C^{3}(\overline{\Omega})}$ and $\partial\Omega$ satisfying
+Then there exists a positive constant $C$ depending only on $\theta_{0}$, $\|\psi\|_{C^{1}(\overline{\Omega}\times[-\mu_{0},\mu_{0}])}$, $\|u\|_{C^{1}(\overline{\Omega})}$ and $\partial\Omega$ satisfying
 (4.1) $\max_{\partial\Omega}|D^{2}u|\leqslant C,$
@@
 Furthermore, we may suppose that in a neighbourhood of the origin, the boundary $\partial\Omega$ is given by
-(4.2) $x_{n}=\rho(x^{\prime})=\frac{1}{2}\sum_{\alpha<n}\kappa^{b}_{\alpha}x_{\alpha}^{2}+O(|x^{\prime}|^{3}),$
-where $\kappa^{b}_{1},\ldots,\kappa^{b}_{n-1}$ are the principal curvatures of $\partial\Omega$ at the origin and $x^{\prime}=(x_{1},\ldots,x_{n-1})$. Define
+(4.2) $x_{n}=\rho(x^{\prime})=\frac{1}{2}\sum_{\alpha<n}\kappa^{\prime}_{\alpha}x_{\alpha}^{2}+O(|x^{\prime}|^{3}),$
+where $\kappa^{\prime}_{1},\ldots,\kappa^{\prime}_{n-1}$ are the principal curvatures of $\partial\Omega$ at the origin as before and $x^{\prime}=(x_{1},\ldots,x_{n-1})$. Define
 $\omega_{\delta}=\{x\in\Omega:\rho(x^{\prime})<x_{n}<\rho(x^{\prime})+\delta^{2},|x^{\prime}|<\delta\},$
@@
 Part 1. The double tangential derivative estimate on $\partial\Omega$, that is
-(4.3) $|u_{\alpha\beta}(0)|\leqslant C\mbox{ for }1\leqslant\alpha,\beta\leqslant n-1,$
+(4.3) $|u_{\alpha\beta}(0)|\leqslant C\mbox{ for }1\leqslant\alpha,\beta\leqslant n-1.$
 Since (4.2), We can obtain (4.3) by differentiating the boundary condition $u=0$ twice on $\partial\Omega$.
@@
 Assume (4.8) is correct, next, we establish the mixed tangential-normal derivative estimate (4.4).
-Since $\Omega$ is $(\eta,k)$-convex, there exist two positive constants $\theta$ and $K$ satisfying
-(4.9) $(\kappa_{1}^{b}-3\theta,\ldots,\kappa_{n-1}^{b}-3\theta,2K)\in\widetilde{\Gamma}_{k}.$
+Note that there exist positive constants $\theta$ and $K$ such that
+(4.9) $(\kappa_{1}^{\prime}-3\theta,\ldots,\kappa_{n-1}^{\prime}-3\theta,2K)\in\widetilde{\Gamma}_{k}$
+holds.
 Define
 (4.10) $v=\rho(x^{\prime})-x_{n}-\theta|x^{\prime}|^{2}+Kx_{n}^{2}$
-on $\overline{\omega}_{\delta}$. We see that when $\delta$ depending on $\theta$ and $K$ is sufficiently small, we have
-[…]
-[…]
-[…]
-In view of (4.2) and (4.9), $\lambda(D^{2}v)\in\widetilde{\Gamma}_{k}$ on $\overline{\omega}_{\delta}$.
+on $\overline{\omega}_{\delta}$. We see that when $\delta(\theta,K)$ is sufficiently small, we have
+(4.11) $v\leqslant$ $-\frac{\theta}{2}|x^{\prime}|^{2},$ $\mbox{ on }\partial_{1}\omega_{\delta},$
+$v\leqslant$ $-\frac{\delta^{2}}{2},$ $\mbox{ on }\partial_{2}\omega_{\delta},$
+$v\leqslant$ $-\frac{\theta\delta^{2}}{2},$ $\mbox{ on }\partial_{3}\omega_{\delta}.$
+Besides, by (4.2) and (4.9), we have
+(4.12) $\lambda(D^{2}v)=(\kappa_{1}^{\prime}-2\theta,\ldots,\kappa_{n-1}^{\prime}-2\theta,2K)\in\widetilde{\Gamma}_{k}$
+on $\overline{\omega}_{\delta}$, where $\lambda(D^{2}v)$ denotes the eigenvalues of $D^{2}v$.
 Thus, there exists an uniform constant $\eta_{0}>0$ depending only on $\theta$, $\partial\Omega$ and $K$ satisfying
@@
 (4.13) $\lambda\left(\frac{1}{w}\{\gamma^{is}(v_{st}-2\eta_{0}\delta_{st})\gamma^{jt}\}\right)\in\widetilde{\Gamma}_{k}\mbox{ on }\overline{\omega}_{\delta}.$
-Actually, we can assume $D^{2}v=\operatorname{diag}(\kappa_{1}^{b}-2\theta,\dots,\kappa_{n-1}^{b}-2\theta,2K)$ by rotating the coordinate system if necessary and note that $\gamma^{ij}=\operatorname{diag}(1,\dots,1,\frac{1}{w})$ at 00 since $|Du(0)|=u_{n}(0)$.
+Actually, we can assume $D^{2}v=\operatorname{diag}(\kappa_{1}^{\prime}-2\theta,\dots,\kappa_{n-1}^{\prime}-2\theta,2K)$ by rotating the coordinate system if necessary and note that $\gamma^{ij}=\operatorname{diag}(1,\dots,1,\frac{1}{w})$ at 0 since $|Du(0)|=u_{n}(0)$.
 It follows that
-$\{\frac{1}{w}(\gamma^{ik}D_{kl}v\gamma^{lj})\}=\operatorname{diag}(\frac{\kappa_{1}^{b}-2\theta}{w},\dots,\frac{\kappa_{n-1}^{b}-2\theta}{w},\frac{2K}{w^{3}}).$
-Let $\kappa_{v}=(\frac{\kappa_{1}^{b}-2\theta}{w},\dots,\frac{\kappa_{n-1}^{b}-2\theta}{w},\frac{2K}{w^{3}})=(\kappa_{v}^{\prime},\frac{2K}{w^{3}})$, we have
-$\frac{\sigma_{k}}{\sigma_{l}}(\kappa_{v})>\frac{\sigma_{k}}{\sigma_{l}}(\kappa_{v}^{\prime},\frac{2K}{w})>0$
-by $w<1$, (4.9) and (2.6). Then we can find a suitable constant $\delta$ such that (4.12) holds.
+$\left\{\frac{1}{w}(\gamma^{ik}D_{kl}v\gamma^{lj})\right\}=\operatorname{diag}\left(\frac{\kappa_{1}^{\prime}-2\theta}{w},\dots,\frac{\kappa_{n-1}^{\prime}-2\theta}{w},\frac{2K}{w^{3}}\right).$
+Let $\kappa_{v}=(\frac{\kappa_{1}^{\prime}-2\theta}{w},\dots,\frac{\kappa_{n-1}^{\prime}-2\theta}{w},\frac{2K}{w^{3}})=(\kappa_{v}^{\prime},\frac{2K}{w^{3}})$.
+By (4.12) and $w<1$, we have $(\kappa_{v}^{\prime},\frac{2K}{w})\in\widetilde{\Gamma}_{k}$.
+Then it follows from the ellipticity of $\sigma_{k}$ in $\widetilde{\Gamma}_{k}$ that $\kappa_{v}\in\widetilde{\Gamma}_{k}$.
+Hence we can find a suitable constant $\delta$ such that (4.13) holds.
 As [22] and [21], we consider the following barrier on $\overline{\omega}_{\delta}$, for sufficiently small $\delta$,
@@
 where $v(x)$ is defined by (4.10), $d(x):=\mathrm{dist}(x,\partial\Omega)$ is the distance from $x$ to the boundary $\partial\Omega$, and $t,N$ are two positive constants to be determined later.
-Since $f^{1/(k-l)}$ is concave in $\Gamma$ and homogeneous of degree one and $|Dd|\equiv 1$ on the boundary $\partial\Omega$, by (4.12) and (2.3), we have,
+Since $f^{1/(k-l)}$ is concave in $\widetilde{\Gamma}_{k}$ and homogeneous of degree one and $|Dd|\equiv 1$ on the boundary $\partial\Omega$, by (4.13) and (2.4), we have,
 $\frac{1}{k-1}\psi^{\frac{1}{k-l}-1}G^{ij}(D^{2}v-\eta_{0}I+NDd\otimes Dd)_{ij}$
@@
 where $R$ is a large undetermined positive constant. We shall prove
-(4.17) Φ0 on $\Phi\leqslant 0\mbox{ on }\overline{\omega}_{\delta}$
-by choosing suitable positive constants $\delta$, $t$, $N$ and $R$.
+(4.18) $\Phi\leqslant 0\mbox{ on }\overline{\omega}_{\delta}$
+by choosing suitable positive constants $\delta$, $t$, NN and $R$.
 We first consider the case that the maximum of $\Phi$ is achieved at an interior point $x_{0}\in\omega_{\delta}$. It follows that at $x_{0}$,
@@
 We may further assume $\delta<2t/N$ so that
-(4.19) $-td+\frac{N}{2}d^{2}\leqslant 0\mbox{ on }\overline{\omega}_{\delta}.$
+(4.20) $td+\frac{N}{2}d^{2}\leqslant 0\mbox{ on }\overline{\omega}_{\delta}.$
 By (4.11) and (4.20), we can conclude that $\Phi\leqslant 0$ on $\partial\omega_{\delta}$ by choosing $R$ larger and then (4.18) is proved.
@@
 at the origin, we have
-$a_{ij}=\begin{pmatrix}\frac{u_{11}}{\omega}&0&\cdots&0&\frac{u_{1n}}{\omega}\\ 0&\frac{u_{22}}{\omega}&\cdots&0&\frac{u_{2n}}{\omega}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\frac{u_{n-1,n-1}}{\omega}&\frac{u_{n-1,n}}{\omega}\\ \frac{u_{n1}}{\omega}&\frac{u_{n2}}{\omega}&\cdots&\frac{u_{n,n-1}}{\omega}&\frac{u_{nn}}{\omega^{3}}\end{pmatrix}$
+$a_{ij}=\begin{pmatrix}\frac{u_{11}}{w}&0&\cdots&0&\frac{u_{1n}}{w}\\ 0&\frac{u_{22}}{w}&\cdots&0&\frac{u_{2n}}{w}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\frac{u_{n-1,n-1}}{w}&\frac{u_{n-1,n}}{w}\\ \frac{u_{n1}}{w}&\frac{u_{n2}}{w}&\cdots&\frac{u_{n,n-1}}{w}&\frac{u_{nn}}{w^{3}}\end{pmatrix}$
 By Lemma 1.2 of [4] and the estimates of $u_{\alpha\beta}(0)$, $u_{\alpha n}(0)$, there exists a constant $R_{1}>0$ sufficiently large such that if $u_{nn}>R_{1}$, then
-$\begin{cases}\kappa_{\alpha}=\dfrac{u_{\alpha\alpha}}{\omega}+o(1),&\alpha=1,\cdots,n-1,\\[10.0pt] \kappa_{n}=\dfrac{u_{nn}}{\omega^{3}}+O(1),\end{cases}$
+$\begin{cases}\kappa_{\alpha}=\dfrac{u_{\alpha\alpha}}{w}+o(1),&\alpha=1,\cdots,n-1,\\[10.0pt] \kappa_{n}=\dfrac{u_{nn}}{w^{3}}+O(1),\end{cases}$
 where $\kappa_{1},\ldots,\kappa_{n}$ are the eigenvalues of $\{a_{ij}\}$.
-Since k ¡ n, by choosing $R_{1}$ large enough, we have at the origin,
-$\psi=\Big{(}\frac{\sigma_{k}}{\sigma_{l}}\Big{)}(\lambda(\eta))=\frac{\big{(}C_{n}^{k}-(n-k+1)\big{)}u_{nn}^{k}/(w^{3k})+O(u_{nn}^{k-1})}{(C_{n}^{l}-(n-l+1))u_{nn}^{l}/(w^{3l})+O(u_{nn}^{l-1})}.$
+Since $k<n$, by choosing $R_{1}$ large enough, we have at the origin,
+$\psi=\Big(\frac{\sigma_{k}}{\sigma_{l}}\Big)(\lambda(\eta))=\frac{C_{n-1}^{k}u_{nn}^{k}/(w^{3k})+O(u_{nn}^{k-1})}{C_{n-1}^{l}u_{nn}^{l}/(w^{3l})+O(u_{nn}^{l-1})}.$
 It then implies the uniform upper bound of $|u_{nn}(0)|$. Combining (4.3) and (4.4), we can obtain the uniform upper bound of $|u_{ij}(0)|$. Then (4.21) follows immediately and Theorem 4.1 is proved.
@@
 Then
-$-G^{s}W_{s}\leqslant C|DW|+\frac{\epsilon_{0}}{2}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon_{0}}G^{ij}W_{i}W_{j}.$
+$G^{s}W_{s}\leqslant C|DW|+\frac{\epsilon_{0}}{2}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon_{0}}G^{ij}W_{i}W_{j}.$
 Obviously, for any $\epsilon_{1}>0$,
@@
 $\gamma^{sj}b_{jr}W_{s}$ $=\gamma^{sj}b_{jr}(u_{\alpha s}+\rho_{\alpha s}u_{n}+\rho_{\alpha}u_{ns}-\sum_{\beta\leqslant n-1}u_{\beta}u_{\beta s})$
-$=\omega\kappa_{r}(\eta_{\alpha r}+\rho_{\alpha}\eta_{nr}-\sum_{\beta\leqslant n-1}u_{\beta}\eta_{\beta r})+\gamma^{sj}b_{jr}\rho_{\alpha s}u_{n}.$
+$=w\kappa_{r}(\eta_{\alpha r}+\rho_{\alpha}\eta_{nr}-\sum_{\beta\leqslant n-1}u_{\beta}\eta_{\beta r})+\gamma^{sj}b_{jr}\rho_{\alpha s}u_{n}.$
 It follows that
-(4.35) $|\gamma^{sj}b_{jr}W_{s}|\leqslant C\omega\kappa_{r}(\epsilon+|\rho_{\alpha}|)+C.$
+(4.36) $|\gamma^{sj}b_{jr}W_{s}|\leqslant Cw\kappa_{r}(\epsilon+|\rho_{\alpha}|)+C.$
 It is obvious that $f_{r}\kappa_{r}=(k-l)\psi-\sum_{i\neq r}f_{i}\kappa_{i}$, thus by $\kappa_{n}^{2}>\epsilon_{0}^{2}\kappa_{r}^{2}$,
-$\frac{2}{\omega}f_{r}\kappa_{r}|(\sum_{t}b_{tr}u_{t})\gamma^{sj}b_{jr}W_{s}|=$ $\frac{2}{\omega}((k-l)\psi-\sum_{i\neq r}f_{i}\kappa_{i})\left|\left(\sum_{t}b_{tr}u_{t}\right)\gamma^{sj}b_{jr}W_{s}\right|$
+$\frac{2}{w}f_{r}\kappa_{r}|(\sum_{t}b_{tr}u_{t})\gamma^{sj}b_{jr}W_{s}|=$ $\frac{2}{w}((k-l)\psi-\sum_{i\neq r}f_{i}\kappa_{i})\left|\left(\sum_{t}b_{tr}u_{t}\right)\gamma^{sj}b_{jr}W_{s}\right|$
 $\leqslant$ $C|DW|+C(\epsilon+|\rho_{\alpha}|)\sum_{i\neq r}f_{i}|\kappa_{i}|\kappa_{r}+C\sum_{i\neq r}f_{i}|\kappa_{i}|$
@@
 $+C\epsilon_{0}(\epsilon+|\rho_{\alpha}|)\sum_{i\neq r}f_{i}\kappa_{r}^{2}+C\epsilon_{1}\sum_{i\neq r}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon_{1}}\sum_{i\neq r}f_{i}$
-$\leqslant$ $C\psi|DW|+(C\epsilon_{0}^{-1}(\epsilon+|\rho_{\alpha}|)+\epsilon_{1})\sum_{i\neq r}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon_{1}}\sum_{i}G^{ii}$
+$\leqslant$ $C|DW|+(C\epsilon_{0}^{-1}(\epsilon+|\rho_{\alpha}|)+\epsilon_{1})\sum_{i\neq r}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon_{1}}\sum_{i}G^{ii}$
 for any $\epsilon_{1}>0$.
@@
 $\left(\epsilon_{0}^{-1}C(\epsilon+|\rho_{\alpha}|)+\epsilon_{1}\right)<\frac{mc_{1}}{4},$
-where $m=1-\theta^{2}$, $\theta$ is defined in spacelike. Therefore, as in Case (b-1), (4.8) follows by (4.27). ∎
+where $m=1-\theta_{0}^{2}$, $\theta_{0}$ is defined in spacelike. Therefore, as in Case (b-1), (4.8) follows by (4.28). ∎
 The last step is to estimate the second derivatives in the interior.
@@
 The second fundamental form $h_{ab}$ satisfies
-(4.36) $F^{ij}\nabla_{i}\nabla_{j}h_{ab}=$ $-F^{ij,kl}\nabla_{a}h_{ij}\nabla_{b}h_{kl}-F^{ij}h_{ij}h_{ak}h_{bk}$
-$+F^{ij}h_{ik}h_{jk}h_{ab}+\nabla_{a}\nabla_{b}\psi$
+(4.37) $F^{ij}\nabla_{ij}h_{ab}=$ $-F^{ij,kl}\nabla_{a}h_{ij}\nabla_{b}h_{kl}-F^{ij}h_{ij}h_{ak}h_{bk}$
+$+F^{ij}h_{ik}h_{jk}h_{ab}+\nabla_{ab}\psi$
 where
@@
 Using (4.37) and differentiating the equation (1.1) twice, we have
-(4.37) $0\geqslant\tilde{F}^{ii}\nabla_{ii}h_{11}=-\tilde{F}^{ij,kl}\nabla_{1}h_{ij}\nabla_{1}h_{kl}-\tilde{\psi}h_{11}^{2}+h_{11}\tilde{F}^{ii}h_{ii}^{2}+\nabla_{11}(\tilde{\psi}).$
+[…]
 where $\tilde{\psi}=\psi^{1/(k-l)}$ and
@@
 (4.39) $\max_{\overline{\Omega}}|D^{2}u|\leqslant C.$
-Proof of theorem 1.3.
-Combined (3.2), (3.1), (4.38), we can obtain
-(4.39) $||u||_{C^{2}(\overline{\Omega})}\leqslant C.$
-Applying the Evans-Krylov theory, we can obtain the $C^{2,\alpha}$ estimates for $u$.
-Using the continuity method and the maximum principle, we can finally establish the existence and uniqueness of solutions to (1.3).
-∎
+We have done.
+5. Proof of the main result
+The proof of Theorem 1.3 relies on a standard continuity method [5]. Let
+$\psi^{0}(x,\underline{u})=\frac{\sigma_{k}}{\sigma_{l}}\left(\lambda(\eta[M_{\underline{u}}])\right)$
+and extend $\psi^{0}$ to $\Omega\times\mathbb{R}$ with $\psi^{0}(x,t)=\psi^{0}(x,\underline{u})$. Extend $\psi$ in the same manner, we consider, for each $t\in[0,1]$,
+(5.1) $\left\{\begin{aligned} &\frac{\sigma_{k}}{\sigma_{l}}\left(\lambda(\eta[M_{u^{t}}])\right)=t\psi+(1-t)\psi^{0}&&\mbox{in}~\Omega,\\ &u^{t}=0&&\mbox{on}~\partial\Omega.\end{aligned}\right$.
+The set $S$ consists of numbers $t\in[0,1]$ such that there exists a function $u^{t}\in C^{2,\alpha}(\overline{\Omega})$, which is an admissible solution of (5.1).
+Obviously, $\underline{u}$ is the solution of (5.1) for $t=0$, so the set $S$ is nonempty.
+We shall prove that $S$ is both open and close, so $S=[0,1]$ and the proof is finished.
+First, by the implict function theorem, we can obtain $S$ is an open set.
+Next, let $\{s_{i}\}_{i\in\mathbb{N}}$ be an arbitrary sequence from the set $S$ and $s\in[0,1]$ satisfies $\lim_{i\to\infty}s_{i}=s$, we shall prove $s\in S$.
+Combine with the estimates
+$\sup_{\overline{\Omega}}|Du^{t}|\leqslant 1-\theta_{0}<1$
+and
+$\|u^{t}\|_{C^{2}(\overline{\Omega})}\leqslant C,$
+we obtain the uniform ellipticity of $\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta))$.
+Since $(\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta)))^{\frac{1}{k-l}}$ is concave (see [28]) in $\widetilde{\Gamma}_{k}$, then by Evans–Krylov theorem, we can obtain a $C^{2,\alpha}$ estimate, which means
+$\|u^{s_{i}}\|_{C^{2,\alpha}(\overline{\Omega})}\leqslant C,$
+where $C$ is independent on $s_{i}$. By the Arzela-Ascoli theorem, we have $u^{s}\in C^{2,\delta}(\overline{\Omega})$ for $\delta\in(0,\alpha)$. Therefore, $s\in S$ and $S$ is closed.
+In addition, by standard nonlinear elliptic regularity, we can obtain $u\in C^{\infty}(\overline{\Omega})$.
+The uniqueness of the solution follows from the maximum principle.
+Acknowledgment. We would like to thank Professor Heming Jiao and the referees for helpful comments and valuable suggestions.